### v1

Abstract
We consider a mean-field optimal control problem with general dynamics including common noise and jumps and show that its minimizers are Nash equilibria of an associated mean-field game of controls. These types of games are necessarily potential, and the Nash equilibria obtained as the minimizers of the control problems are naturally related to the McKean-Vlasov equations of Langevin type. We provide several examples to motivate the general theory including the Cucker-Smale Flocking and Kuramoto mean-field games. The invariance property of the value function of these control problems, utilized in our proof, is also proved.
1 Introduction
Mean-field games (MFGs) were independently introduced by Lasry & Lions [46, 47, 48] and Huang, Caines, & Malhamé [35, 36, 37, 38] to approximate large population games in which the interaction appears through the empirical distribution of all agents. Since their introduction, they have been widely applied in contexts ranging from heterogeneous agents models in economics to models of jet-lag recovery in neuroscience.
We study a special class of MFGs that arise as first-order conditions of a mean-field control problem (MFC). This connection has first been observed in the seminal work [48] by Lasry & Lions and has been subsequently well-studied under the name potential mean-field games as we review below. Our primary aim is to explore the natural links between the potential MFG and MFC problems as well as the links with a class of Wasserstein gradient flows. The latter connection can be leveraged to construct compelling mean-field games that have similar qualitative behavior as the gradient flows, and in Section 6 we discuss two such examples: the MFG flocking model of [54, 55] and the Kuramoto MFG of [14, 60, 61].
Specifically, let ${\mathscr{P}}_{2}(\mathbb{R}^{n})$ be the Wasserstein space of probability measures on $\mathbb{R}^{n}$ with finite second moments. For a sufficiently differentiable function $F:{\mathscr{P}}_{2}(\mathbb{R}^{n})\to\mathbb{R}$, consider the following McKean-Vlasov stochastic differential equation:
$\mathrm{d}X_{t}=-\nabla_{x}\delta_{\nu}F({\mathcal{L}}(X_{t}))(X_{t})\,\mathrm{d}t+\sigma\mathrm{d}W_{t},$ (1.1)
where $\delta_{\mu}F(\mu)(x)$ is the linear derivative of $F$ at $\mu\in{\mathscr{P}}_{2}(\mathbb{R}^{n})$, ${\mathcal{L}}(X_{t})$ denotes the law of $X_{t}$, and $W$ is a standard Brownian motion in $\mathbb{R}^{n}$. In analogy to finite-dimensional Langevin dynamical systems, [34] refers to (1.1) as the mean-field Langevin dynamics.
Recently, these equations have received considerable attention due to their connection to noisy gradient descent algorithms in deep neural networks [19, 34, 52, 53]. Additionally, we argue in Section 6 that one can construct potential MFC problems with similar qualitative behavior. In particular, the optimally controlled state dynamics are in the above form (1.1).
While most existing studies on potential games, such as [4], start from a MFG and establish structural conditions on the given functions under which equilibria can be obtained from a MFC problem, we start from a general MFC problem and construct an associated potential MFG. In summary, we propose the following relations,
MF-Langevin Dynamics MFC Problem Potential MFG
Additionally, by working with the probabilistic formulation of both the control and game problems, we obtain general results with non-Markovian controlled dynamics under jumps and common noise. Our approach extends to cost functions that depend on the joint distribution of the state and control and which are non-separable in the control and measure variable. This further allows us to derive mean-field games of controls.
MFC, MFG, and Main Results.
On a filtered probability space $(\Omega,{\mathscr{F}},\mathbb{F},\mathbb{P})$ satisfying the usual conditions, let $W^{0}$ and $W$ be $\mathbb{F}$-Brownian motions, and $N$ and $N^{0}$ be stationary $\mathbb{F}$-Poisson random measures on $[0,\infty)\times(\mathbb{R}^{k}\setminus\{0\})$. In our setting, $W$ and $N$ represent the idiosyncratic noise, while $W^{0}$ and $N^{0}$ model the common noise. Let $T>0$ be the time horizon. For a given initial condition $X_{0}$, and coefficients $(b,\sigma,\sigma^{0},\lambda,\lambda^{0})$ satisfying Assumption 2.3 below, the controlled state process $X^{\alpha}\in\mathbb{R}^{n}$ is the unique strong solution to the following stochastic differential equation,
$X^{\alpha}_{t}=$ $X_{0}+\int_{0}^{t}b(s,X^{\alpha}_{\cdot\land s},\alpha_{s})\,\mathrm{d}s+\int_{0}^{t}\left(\sigma(s,X^{\alpha}_{\cdot\land s},\alpha_{s})\,\mathrm{d}W_{s}+\sigma^{0}(s,X^{\alpha}_{\cdot\land s},\alpha_{s})\,\mathrm{d}W^{0}_{s}\right)$ (1.2)
$+\int_{0}^{t}\int_{\mathbb{R}^{k}\setminus\{0\}}\left(\lambda(s,X^{\alpha}_{\cdot\land s-},\alpha_{s},\zeta)\,\widetilde{N}(\mathrm{d}s,\mathrm{d}\zeta)+\lambda^{0}(s,X^{\alpha}_{\cdot\land s-},\alpha_{s},\zeta)\,\ \widetilde{N}^{0}(\mathrm{d}s,\mathrm{d}\zeta)\right),$
where $\alpha_{\cdot}$ is any $\mathbb{F}$-predictable, $\mathrm{d}\mathbb{P}\otimes\mathrm{d}t$-square integrable control process taking values in a Borel subset $U$ of some Euclidean space. Let ${\mathscr{A}}$ denote the set of all such control processes. Notice that the above dynamics do not contain any mean field interactions, and this is our main structural assumption.
Consider the cost functionals,
$F:[0,T]\times{\mathscr{D}}_{n}\times U\times{\mathscr{P}}({\mathscr{D}}_{n}\times U)\mapsto\mathbb{R},\quad G:{\mathscr{D}}_{n}\times{\mathscr{P}}({\mathscr{D}}_{n})\mapsto\mathbb{R},$
where ${\mathscr{D}}_{n}$ is the Skorokhod space of càdlàg functions from $[0,T]$ to $\mathbb{R}^{n}$, and for any measurable space $S$, ${\mathscr{P}}(S)$ is the set of probability measures on $S$. We set
$J(X_{0},\alpha):=\mathbb{E}\left[\int_{0}^{T}F(t,X^{\alpha}_{\cdot\wedge t},\alpha_{t},{\mathcal{L}}^{\alpha}_{t})\,\mathrm{d}t+G(X^{\alpha}_{\cdot\land T},{\mathcal{L}}^{\alpha}_{T})\ \right],$ (1.3)
where ${\mathcal{L}}^{\alpha}_{t}:={\mathcal{L}}((X^{\alpha}_{\cdot\land t},\alpha_{t})\,|\,W^{0},N^{0})$ is the regular conditional distribution of $(X^{\alpha}_{\cdot\land t},\alpha_{t})$ given the common noise $W^{0},N^{0}$. Then, the mean-field control (MFC) problem consists of minimizing $J(X_{0},\alpha)$ over all $\alpha\in{\mathscr{A}}$.
We next introduce the associated mean-field game of controls. Let $\Pi:{\mathscr{D}}_{n}\times U\mapsto{\mathscr{D}}_{n}$ be the projection on the first coordinate, and for $(t,x,a,\nu)\in[0,T]\times{\mathscr{D}}_{n}\times U\times{\mathscr{P}}({\mathscr{D}}_{n}\times U)$, let $\mu=\nu\circ\Pi^{-1}$ and set
$\begin{split}f(t,x,a,\nu)&:=\int_{{\mathscr{D}}_{n}\times U}\delta_{\nu}F(t,\tilde{x},\tilde{a},\nu)(x,a)\,\nu(\mathrm{d}\tilde{x},\mathrm{d}\tilde{a})+F(t,x,a,\nu),\\ g(x,\nu)&:=\int_{{\mathscr{D}}_{n}}\delta_{\mu}G(\tilde{x},\mu)(x)\,\mu(\mathrm{d}\tilde{x})+G(x,\mu).\end{split}$ (1.4)
Given an adapted stochastic process $\nu^{*}=(\nu^{*}_{t})_{t\in[0,T]}$ with values in ${\mathscr{P}}({\mathscr{D}}_{n}\times U)$, we set
$J_{g}(X_{0},\alpha,\nu^{*}):=\mathbb{E}\left[\int_{0}^{T}f(t,X^{\alpha}_{\cdot\wedge t},\alpha_{t},\nu^{*}_{t})\,\mathrm{d}t+g(X^{\alpha}_{\cdot\land T},\nu^{*}_{T})\right].$
Then, a Nash equilibrium of the associated MFG of controls is a pair of an admissible control process $\alpha^{*}\in{\mathscr{A}}$ and a process $\nu^{*}=(\nu^{*}_{t})_{t\in[0,T]}$ with values in ${\mathscr{P}}({\mathscr{D}}_{n}\times U)$ which satisfy,
$J_{g}(X_{0},\alpha^{*},\nu^{*})=\inf_{\alpha\in{\mathscr{A}}}\,\,J_{g}(X_{0},\alpha,\nu^{*})\ \ \text{and}\ \ \nu^{*}_{t}={\mathcal{L}}((X^{\alpha^{*}}_{\cdot\land t},\alpha^{*}_{t})\,|\,W^{0},N^{0}),\ \ \forall\ t\in[0,T].$ (1.5)
The main result of the paper connects the MFC and MFG problems.
Theorem.
Under natural regularity and growth conditions on the coefficients, for any minimizer $\alpha^{*}$ of the MFC problem (1.3), the pair $(\alpha^{*},\nu^{*})$ with $\nu^{*}_{t}:={\mathcal{L}}((X^{\alpha^{*}}_{\cdot\land t},\alpha^{*}_{t})\,|\,W^{0},N^{0})$ for $t\in[0,T]$, is a Nash equilibrium of the MFG of controls (1.5).
The precise formulation is given in Theorem 3.6 below. Additionally, in Section 3 and subsection 4.3 below, we discuss specific examples and structures, and show that this theorem covers all previous results. In particular, it is clear that the results of this paper remain true in the absence of the common noise by simply setting $\sigma^{0}=\lambda^{0}=0$ and in the absence of jumps by setting $\lambda=\lambda^{0}=0$. Moreover, as discussed in subsection 4.2, in the classical separable models, we obtain a standard MFG depending only on the distribution of the state process and not the control.
One technical tool we utilize is the invariance principle Theorem 7.2. It states that the value function of MFC and the Nash equilibria of MFG do not depend on the underlying probabilistic structure. This property is needed in our approach since in our proof we expand the initially given probabilistic structure by adding independent Bernoulli random variables. For diffusions it is established in [20, 25], and we extend it to our setting.
Context and Earlier Results.
Succeeding the introduction of potential MFGs in [48, Section 2.6], they have been studied in [4, 6, 7, 10, 56] in both the first-order (i.e. without Brownian noise) and the second-order cases. The majority of the existing literature on potential mean-field games starts from a characterization of the Nash equilibria as solutions of a forward-backward system of partial differential equations. These works interpret the PDE system as an optimality condition for two problems in duality by using an elegant application of the Fenchel-Rockafellar Duality Theorem or the von Neumann Minimax Principle, as illustrated in [56]. For the first-order potential MFGs we refer readers to the recent paper [30] which provides an excellent introduction to these techniques and also a historical context. A similar analysis in the context of optimal transport can be found in [9] and for (first-order) variational MFGs in [1, Section 2]. Extensions of these results include infinite-horizon cost criteria [51] and problems with state constraints [22]. The structure of potential MFGs has been leveraged to study weak solutions of the master equation [16] and to prove the convergence of learning procedures to solutions of the MFG system [4, 8]. We also refer to [32] for the study of stochastic differential potential games with finitely many players.
On finite state spaces, direct arguments such as the Pontryagin principle imply that potential MFG equilibria can be viewed as first-order conditions of an associated control problem as initiated in [29, 31]. For MFGs with finite state space, the long-time behavior is investigated in [27]; the convergence problem of the $N$-player game in a two-state model without uniqueness is examined in [17]. For finite state problems with no uniqueness, [15] proposes a selection principle of MFG equilibria as the minimizers of the associated control problem when uniqueness fails. In the continuous state case, an approach to potential MFGs that is not based on convex duality is presented in the classical book [11]. Chapters 6.2.5 and 6.7.2 of [11] identify the Pontryagin system of the MFC problem with the FBSDE system describing MFG Nash equilibria in the probabilistic setting.
We work in a general stochastic formulation of the control problem using direct methods, rather than the FBSDEs. Our results are in considerable generality, extending to $(i)$ the environments with common noise, $(ii)$ the models with jumps, $(iii)$ the cost functionals $F,G$ that depend on the control and measure components in a non-separable way, and $(iv)$ coefficients that depend on the entire path of the state variable. The non-separability of the cost functions leads to a mean-field game of controls, where the interaction between agents is through both the state and control variables. Further discussion is given in subsection 4.3.
The paper is organized as follows: After reviewing the notation and the setting, in Section 3 we define the MFC problem and the MFG of controls, and then state the main result. We discuss the result and several special cases in Section 4, and the proof of our main result Theorem 3.6 is given in Section 5. We propose a connection between the Langevin dynamics, MFC, and MFG, and illustrate it through two examples in Section 6. Section 7 proves the invariance principle. Appendix A provides a summary of standard results and notations on point processes, and Appendix B discusses the discounted infinite horizon setting.
2 Notation and Setting
For a topological space $S$, ${\mathscr{B}}(S)$ is the set of all Borel sets. The dimension of the state process is denoted by $n$, $d$ refers to the dimension of the idiosyncratic Brownian motion, and $\ell$ is the dimension of the common Brownian noise. $T>0$ is the horizon. ${\mathscr{D}}_{k}(0,T)$ is the Polish space of càdlàg functions from $[0,T]$ to $\mathbb{R}^{k}$ endowed it with the Skorokhod $J_{1}$ topology. When the domain is clear, we write ${\mathscr{D}}_{k}$. On ${\mathscr{D}}_{k}$, we sometimes utilize the supremum norm $\|x\|_{T}:=\sup_{t\in[0,T]}|x(t)|$. The control set $U$ is a non-empty Borel subset of a Euclidean space. On any Polish space $S$, ${\mathscr{P}}(S)$ is the set of probability measures, and ${\mathscr{P}}_{2}(S)$ is the Wasserstein space of measures with finite second moments. A function $\phi:{\mathscr{P}}_{2}(S)\to\mathbb{R}$ is called (linearly) differentiable, if there exists a measurable function $\delta_{\eta}\phi:{\mathscr{P}}_{2}(S)\times S\to\mathbb{R}$ which has at most of quadratic growth in $x\in S$ and satisfies,
$\phi(\mu)-\phi(\eta)=\int_{0}^{1}\int_{S}\delta_{\eta}\phi(\eta+\tau(\mu-\eta))(s)\,(\mu-\eta)(\mathrm{d}s)\,\mathrm{d}\tau,\qquad\forall\mu,\eta\in{\mathscr{P}}_{2}(S).$
For a $\sigma$-algebra ${\mathscr{G}}$ and an $S$-valued random variable $Y$, the random measure ${\mathcal{L}}(Y\,|\,{\mathscr{G}})\in{\mathscr{P}}(S)$ is the regular condition distribution of $Y$ given ${\mathscr{G}}$. We write ${\mathcal{L}}^{\mathbb{P}}={\mathcal{L}}$, when the dependence on the underlying probability measure $\mathbb{P}$ is important. When ${\mathscr{G}}$ is trivial, ${\mathcal{L}}(Y)$ is the law of $Y$. Finally, ${\mathscr{L}}^{2}({\mathscr{G}})$ denotes the set of all square-integrable $n$-dimensional random variables which are ${\mathscr{G}}$-measurable.
2.1 Probabilistic Structures
We follow the standard terminology of [39] and provide further details in Appendix A.
On the jump space $(E,{\mathscr{E}}):=(\mathbb{R}^{k}\setminus\{0\},{\mathscr{B}}(\mathbb{R}^{k}\setminus\{0\}))$ we fix characteristic jump measures $n(\mathrm{d}\zeta)$ and $n^{0}(\mathrm{d}\zeta)$ of the idiosyncratic jumps $N$ and common jumps $N^{0}$, respectively.
Definition 2.1.
A reference probability basis $\gamma$ is a tuple $\gamma=(\Omega,{\mathscr{F}},\mathbb{P},\mathbb{F},{\mathscr{W}},{\mathscr{N}})$ given by,
  1. (i)
    $(\Omega,{\mathscr{F}},\mathbb{P},\mathbb{F}=({\mathscr{F}}_{t})_{t\geq 0})$ is a filtered probability space satisfying the usual conditions;
  2. (ii)
    ${\mathscr{W}}=(W^{0},W)$, where $W^{0}$ and $W$ are $\mathbb{F}$-Brownian motions in $\mathbb{R}^{\ell}$ and $\mathbb{R}^{d}$, respectively;
  3. (iii)
    ${\mathscr{N}}=(N^{0},N)$, where $N^{0}$ and $N$ are Poisson random measures associated to stationary $\mathbb{F}$-Poisson stationary point processes with characteristic jump measures $n(\mathrm{d}\zeta)$ and $n^{0}(\mathrm{d}\zeta)$, respectively.
We assume that the corresponding point processes $p^{0}$ and $p$ have $\mathbb{P}$-almost surely disjoint domains, or equivalently their jump times are almost surely disjoint. Finally, let ${\mathscr{G}}$ denote the completion of the sigma-algebra generated by the common noise processes $(W^{0},N^{0})$.
Definition 2.2 (Admissible controls).
The set of admissible controls ${\mathscr{A}}$ is the collection of all $\mathbb{F}$-predictable processes $\alpha:\Omega\times[0,T]\to U$ that are square integrable, i.e.,
$\|\alpha\|_{2,T}^{2}:=\mathbb{E}\int_{0}^{T}|\alpha_{t}|^{2}\ \mathrm{d}t<\infty.$ (2.1)
When the dependence on the basis $\gamma$ is important, we write ${\mathscr{A}}(\gamma)$, ${\mathcal{G}}(\gamma)$,
$\gamma=(\Omega^{\gamma},{\mathscr{F}}^{\gamma},\mathbb{P}^{\gamma},\mathbb{F}^{\gamma}=({\mathscr{F}}_{t}^{\gamma})_{t\in[0,T]},{\mathscr{W}}^{\gamma},{\mathscr{N}}^{\gamma}),\ \ \mathbb{E}^{\gamma}=\mathbb{E}^{\mathbb{P}^{\gamma}},\ \text{and}\ \ {\mathcal{L}}^{\gamma}={\mathcal{L}}^{\mathbb{P}^{\gamma}},$
and we suppress this dependence in our notation when it is clear or redundant.
2.2 Controlled State Process
In this subsection, we state standard conditions on the coefficients of the equation (1.2):
$(b,\sigma,\sigma_{0}):[0,T]\times{\mathscr{D}}_{n}\times U\mapsto\mathbb{R}^{n}\times{\mathscr{M}}_{n\times d}\times{\mathscr{M}}_{n\times\ell},$
$(\lambda,\lambda^{0}):[0,T]\times{\mathscr{D}}_{n}\times U\times E\mapsto\mathbb{R}^{n}\times\mathbb{R}^{n},$
where ${\mathscr{M}}_{m\times m^{\prime}}$ is the set of $m\times m^{\prime}$ matrices.
Assumption 2.3.
All coefficients are measurable and there exist constants $K,L>0$ and $K(\zeta),L(\zeta)$ such that for any $(t,x,a,\zeta)\in[0,T]\times{\mathscr{D}}_{n}\times U\times E$ we have,
$|(b,\sigma,\sigma^{0})(t,x,a)-(b,\sigma,\sigma^{0})(t,x^{\prime},a)|$ $\leq L\|x-x^{\prime}\|_{t},$
$|(\lambda,\lambda^{0})(t,x,a,\zeta)-(\lambda,\lambda^{0})(t,x^{\prime},a,\zeta)|$ $\leq L(\zeta)\|x-x^{\prime}\|_{t},$
$|(b,\sigma,\sigma^{0})(t,x,a)|^{2}\leq K(1+\|x\|_{t}^{2}+|a|^{2}),\ \ |(\lambda,\lambda^{0})(t,x,a,\zeta)|^{2}\leq K(\zeta)^{2}(1+\|x\|_{t}^{2}+|a|^{2}),$
and the constants $L(\zeta),K(\zeta)$ satisfy $\int_{E}(L(\zeta)^{p}+K(\zeta)^{p})\,(n^{0}+n)(\mathrm{d}\zeta)<\infty$ for any $p\geq 2$.
Set $\boldsymbol{\sigma}:=(\sigma^{0},\sigma)$, and $\boldsymbol{\lambda}:=(\lambda^{0},\lambda)$, and we rewrite (1.2) as,
$X^{\alpha}_{t}=X_{0}+\int_{0}^{t}$ $b(s,X^{\alpha}_{\cdot\land s},\alpha_{s})\,\mathrm{d}s+\int_{0}^{t}\boldsymbol{\sigma}(s,X^{\alpha}_{\cdot\land s},\alpha_{s})\,\mathrm{d}{\mathscr{W}}_{s}$
$+\int_{0}^{t}\int_{E}\boldsymbol{\lambda}(s,X^{\alpha}_{\cdot\land s-},\alpha_{s},\zeta)\,\widetilde{\mathscr{N}}(\mathrm{d}s,\mathrm{d}\zeta),\quad t\geq 0.$ (SDE$(X_{0},\alpha)$)
Under these assumptions, the standard Picard iteration technique implies the following theorem, see [42, Theorem 3.1].
Theorem 2.4.
Under Assumption 2.3, for any probability basis $\gamma$, admissible control $\alpha\in{\mathscr{A}}(\gamma)$, and $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}_{0})$, there exists a unique $\mathbb{F}$-adapted càdlàg process $X^{\alpha}$ that satisfies SDE$(X_{0},\alpha)$. Furthermore, $\mathbb{E}^{\mathbb{P}}[\sup_{0\leq s\leq t}|X_{s}^{\alpha}|^{2}]<\infty$ for all $t\geq 0$.
3 Main Result
In this section, we first define the MFC and MFG problems and then state the main result.
3.1 The Mean-Field Control Problem
Let a probability basis $\gamma$ be given and fix $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}_{0})$. For a finite horizon $T<\infty$, let
$F:[0,T]\times{\mathscr{D}}_{n}\times U\times{\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)\mapsto\mathbb{R},\qquad G:{\mathscr{D}}_{n}\times{\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)\mapsto\mathbb{R},$
be given running and terminal cost functions satisfying Assumption 3.5 below. Let $J(X_{0},\alpha)$ be as in (1.3), and recall that ${\mathscr{G}}$ is the completion of the $\sigma$-algebra generated by the common noise processes $(W^{0},N^{0})$. In the absence of common noise, ${\mathscr{G}}$ is the trivial $\sigma$-algebra. We have the following important technical observation.
Lemma 3.1.
The ${\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)$ measure-valued process
${\mathcal{L}}^{\alpha}_{t}={\mathcal{L}}((X^{\alpha}_{\cdot\land t},\alpha_{t})\,|\,{\mathscr{G}})={\mathcal{L}}((X^{\alpha}_{\cdot\land t},\alpha_{t})\,|\,\sigma(W^{0}_{s},N^{0}_{s}([0,s]\times B):s\leq t,B\in{\mathscr{E}}))$
has an $\mathbb{F}$-progressively measurable modification, and we always consider this version.
Proof.
Equipped with the product of the Skorokhod $J_{1}$ and Euclidean topologies, ${\mathscr{D}}_{n}\times U$ is a separable metric space. Since ${\mathscr{F}}_{0}$ is assumed to contain all $({\mathscr{F}},\mathbb{P})$-null sets, ${\mathcal{L}}^{\alpha}_{t}$ is adapted to the filtration $\mathbb{F}$. Therefore, there exists a $\mathbb{F}$-progressively measurable modification of ${\mathcal{L}}^{\alpha}$. ∎
For a given $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}_{0})$, the value function of the MFC is defined by,
$v(X_{0},\gamma):=\inf_{\alpha\in{\mathscr{A}}}J^{\gamma}(\alpha,X_{0})=\inf_{\alpha\in{\mathscr{A}}(\gamma)}\mathbb{E}^{\gamma}\left[\int_{0}^{T}F(t,X^{\alpha}_{\cdot\wedge t},\alpha_{t},{\mathcal{L}}^{\alpha}_{t})\,\mathrm{d}t+G(X^{\alpha}_{\cdot\land T},{\mathcal{L}}^{\alpha}_{T})\right].$ (3.1)
Further, for $\varrho\in{\mathscr{P}}_{2}(\mathbb{R}^{n})$, let $\Gamma(\varrho)$ be the set of all bases $\gamma$ such that there exists $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}_{0}^{\gamma})$ with distribution $\varrho$, let ${\mathcal{I}}^{\gamma}(\varrho)$ be the set of all such initial conditions, and we set
$V(\varrho):=\inf_{\gamma\in\Gamma(\varrho)}\inf_{X_{0}\in{\mathcal{I}}^{\gamma}(\varrho)}v(X_{0},\gamma)=\inf_{\gamma\in\Gamma(\varrho)}\inf_{X_{0}\in{\mathcal{I}}^{\gamma}(\varrho)}\inf_{\alpha\in{\mathscr{A}}}J^{\gamma}(\alpha,X_{0}).$ (3.2)
Under mild conditions on the coefficients, $v$ depends only on the law of $X_{0}$, and we now formalize this property.
Definition 3.2.
We say that an MFC problem is invariant under the probability bases, if for any $\gamma,\tilde{\gamma}$, $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}^{\gamma}_{0})$, $\tilde{X}_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}^{\tilde{\gamma}}_{0})$, and $\varrho\in{\mathscr{P}}_{2}(\mathbb{R}^{n})$ we have,
${\mathcal{L}}^{\gamma}(X_{0})={\mathcal{L}}^{\tilde{\gamma}}(\tilde{X}_{0})=\varrho\quad\Rightarrow\quad v(X_{0},\gamma)=v(\tilde{X}_{0},\tilde{\gamma})=V(\varrho).$
In the diffusion setting, invariance is proved in [20, Theorem 3.6] for ‘rich enough’ probabilistic structures, see [20, Section 2.1.3], and in [26, Proposition 2.4], and our Theorem 7.2 extends it to more general models. Although almost all problems have this property, some natural continuity and growth assumptions are needed as summarized in Assumption 7.1 and the counter example provided in [20, Section 3] highlights their necessity.
In what follows, we sometimes suppress the dependence on $\gamma$ and write $v(X_{0})$, $J(X_{0},\alpha)$. Additionally, although the terminal cost $G$ depends on ${\mathcal{L}}^{\alpha}_{T}$ only through its projection ${\mathcal{L}}(X^{\alpha}_{\cdot\land T}\,|\,{\mathscr{G}})$, this observation is not relevant in our analysis and we do not emphasize it.
3.2 The Potential Mean-Field Game of Controls
In this subsection, we give the definition of the potential mean-field game of controls equilibrium under common noise; see [13, Definition 2.1] for MFGs under common noise and [23, 24] for general MFC and MFG of controls close to our setting.
In these problems, the representative player takes an $\mathbb{F}$-adapted measure valued process $t\mapsto\nu^{*}_{t}\in{\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)$ as given, and solves the optimal control problem of minimizing,
$J_{g}(X_{0},\alpha,\nu^{*}):=\mathbb{E}\left[\int_{0}^{T}f(t,X^{\alpha}_{\cdot\land t},\alpha_{t},\nu^{*}_{t})\,\mathrm{d}t+g(X^{\alpha}_{\cdot\land T},\nu^{*}_{T})\right]$ (3.3)
over all $\alpha\in{\mathscr{A}}$, where $f$, $g$ are as in (1.4). Then, the Nash equilibrium is defined as follows.
Definition 3.3.
Consider an initial condition $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}_{0})$. A pair $(\alpha^{*},\nu^{*})$ is a (potential) MFG of controls Nash equilibrium if the followings hold:
  1. (i)
    The control process $\alpha^{*}\in{\mathscr{A}}$ satisfies $J_{g}(X_{0},\alpha^{*},\nu^{*})=\inf_{\alpha\in{\mathscr{A}}}\ J_{g}(X_{0},\alpha,\nu^{*})$.
  2. (ii)
    $\nu^{*}=(\nu^{*}_{t})_{t\in[0,T]}$ is an $\mathbb{F}$-progressive, ${\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)$-valued process that satisfies
    $\nu^{*}_{t}={\mathcal{L}}^{\alpha^{*}}_{t}={\mathcal{L}}((X^{\alpha^{*}}_{\cdot\land t},\alpha^{*}_{t})\,|\,{\mathscr{G}}),\qquad\mathbb{P}{\text{-a.s.}}\ \forall\ t\in[0,T].$ (3.4)
Remark 3.4.
In the absence of common noise, $\nu^{*}=(\nu^{*}_{t})_{t\in[0,T]}$ is a deterministic ${\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)$-valued process. Then, the second condition above reduces to $\nu^{*}_{t}={\mathcal{L}}(X^{\alpha^{*}}_{\cdot\land t},\alpha^{*}_{t})$ for all $t\in[0,T]$. In this case, Borel measurability of $t\mapsto{\mathcal{L}}(X^{\alpha^{*}}_{\cdot\land t},\alpha^{*}_{t})$ follows directly from the measurability of $(X^{\alpha^{*}},\alpha^{*})$.
3.3 Main Result
Our main result provides a connection between the MFC and MFG problems. It is the central result of the paper and its proof is given in Section 5 by probabilistic arguments.
To state our assumptions, for any $\nu\in{\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)$, we denote the second moment by
$\mathfrak{s}(\nu):=\int_{{\mathscr{D}}_{n}\times U}(\|x\|^{2}_{T}+|a|^{2})\,\nu(\mathrm{d}x,\mathrm{d}a).$ (3.5)
Assumption 3.5.
We assume that $F$,$G$ are measurable and they satisfy:
  • (Differentiability). The functions $F(t,x,a,\cdot)$ and $G(x,\cdot)$ are linearly differentiable.
  • (Quadratic Growth). For all $(t,x,x^{\prime},\nu,a)\in[0,T]\times{\mathscr{D}}_{n}^{2}\times{\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)\times U$,
    $|F(t,x,a,\nu)|+|G(x,\nu)|+$ $\left|\delta_{\nu}F(t,x,a,\nu)(x^{\prime})\right|+\left|\delta_{\nu}G(x,\nu)(x^{\prime})\right|$
    $\leq C(1+\|x\|_{T}^{2}+|a|^{2}+\mathfrak{s}(\nu)+\|x^{\prime}\|_{T}^{2}).$
We prove the main result under the assumption of invariance proved in Theorem 7.2 under Assumption 7.1. Additionally, its proof reveals that it holds under a weaker version of invariance as discussed in Remark 5.3 below.
Theorem 3.6.
Suppose that Assumptions 2.3 and 3.5 hold, and the MFC problem with payoff functional (3.1) is invariant under probability bases. Then, for $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}_{0})$, $\alpha^{*}\in{\mathscr{A}}$, and $\nu^{*}_{t}={\mathcal{L}}((X^{\alpha^{*}}_{\cdot\land t},\alpha^{*}_{t})\,|\,{\mathscr{G}})$,
$J(X_{0},\alpha^{*})=\inf_{\alpha\in{\mathscr{A}}}\ J(X_{0},\alpha)\quad\Rightarrow\quad J_{g}(X_{0},\alpha^{*},\nu^{*})=\inf_{\alpha\in{\mathscr{A}}}\ J_{g}(X_{0},\alpha,\nu^{*}).$
Consequently, for any minimizer $\alpha^{*}\in{\mathscr{A}}$ of the MFC problem (3.1), the pair $(\alpha^{*},\nu^{*})$ is a Nash equilibrium for the MFG of controls (3.3).
As we discuss in the next section, the above result is an extension of earlier results of [4] and the references therein. We additionally show that when the MFC problem has the widely employed separable structure, the resulting MFG is a standard one which depends only on the distribution of the state process and not the control.
4 Discussion and Special Cases
This section discusses special cases that are frequently used in the literature and assumes the absence of common noise $(W^{0},N^{0})$. This assumption turns $t\mapsto{\mathcal{L}}^{\alpha}_{t}$ into a deterministic flow, hence Nash equilibria only involve deterministic processes $t\mapsto\nu_{t}\in{\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)$.
4.1 Separable Running Cost
We first investigate separable cost functions. Assume that for suitable functions,
$F_{0}:[0,T]\times{\mathscr{D}}_{n}\times{\mathscr{P}}_{2}({\mathscr{D}}_{n})\mapsto\mathbb{R},\ \ F_{1}:[0,T]\times{\mathscr{P}}_{2}({\mathscr{D}}_{n})\mapsto\mathbb{R},\ \ G_{0}:{\mathscr{P}}_{2}({\mathscr{D}}_{n})\mapsto\mathbb{R},$
and for $(t,x,\nu,a)\in[0,T]\times{\mathscr{D}}_{n}\times{\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)\times U$ we have,
$F(t,x,a,\nu)=F_{0}(t,x,a)+F_{1}(t,\mu),\quad G(x,\nu)=G_{0}(\mu),$ (4.1)
where $\mu=\nu\circ\Pi^{-1}$ and, as before, $\Pi:{\mathscr{D}}_{n}\times U\mapsto{\mathscr{D}}_{n}$ is the projection on the first variable. This setting corresponds to a standard mean-field game, and in most applications $F_{0}=\tfrac{1}{2}|a|^{2}$. Then, the costs $f(t,x,\mu,a)$, $g(t,\nu)$ given in (1.4) depend only on $\mu=\nu\circ\Pi^{-1}$ and we have
$f(t,x,\nu,a)=F_{0}(t,x,a)+\delta_{\mu}F_{1}(t,\mu)(x)+F_{1}(t,\mu),\quad g(x,\mu)=\delta_{\mu}G(\mu)(x)+G(\mu).$
For a given flow $t\mapsto\mu_{t}\in{\mathscr{P}}_{2}({\mathscr{D}}_{n})$, the terms $F_{1}(t,\mu_{t})$ and $G(\mu_{T})$ are independent of the control. Therefore, they do not influence the control problems and without loss of generality, we omit them in the definitions of the potential MFG.
The optimization of the representative agent in the MFG is to minimize
$J_{g}(X_{0},\alpha,\mu)=\mathbb{E}\left[\int_{0}^{T}\left[F_{0}(t,X^{\alpha}_{\cdot\land t},\alpha_{t})+\delta_{\mu}F_{1}(t,\mu_{t})(X^{\alpha}_{\cdot\land t})\right]\,\mathrm{d}t+\delta_{\mu}G(\mu_{T})(X^{\alpha}_{T})\right],$ (4.2)
over all $\alpha\in{\mathscr{A}}$, where $X^{\alpha}$ is the controlled state SDE$(X_{0},\alpha)$ with $\sigma^{0}=\lambda^{0}=0$.
The following is a direct consequence of Theorem 3.6 and Theorem 7.2.
Corollary 4.1.
Suppose Assumptions 2.3, 3.5 and 7.1 hold, and assume that $\alpha^{*}\in{\mathscr{A}}$ is a minimizer of the MFC problem (3.1) without common noise and with running cost given by (4.1). Then, with $\mu^{*}_{t}={\mathcal{L}}(X^{\alpha^{*}}_{\cdot\land t})$ the pair $(\alpha^{*},\mu^{*})$ is a Nash equilibrium for the potential MFG with the cost functional (4.2).
4.2 Markovian Cost Without Jumps
In addition to separability of the cost functionals, we now assume that $\lambda=0$. Then, the state process $X^{\alpha}$ is continuous in time, and the invariance is proved in [26, Proposition 2.4]. Following [4], we also assume that the cost functions in (4.1) are Markovian, i.e. at time $t\in[0,T]$ the dependence on the state $X$ is only through the current position $X_{t}$. Hence, the payoff functional (3.1) has the form,
$J(\alpha,X_{0})=\mathbb{E}\left[\int_{0}^{T}\left(\tilde{F}_{0}(t,X^{\alpha}_{t},\alpha_{t})+\tilde{F}_{1}(t,{\mathcal{L}}(X^{\alpha}_{t}))\right)\,\mathrm{d}t\right]+\tilde{G}({\mathcal{L}}(X^{\alpha}_{T})),$ (4.3)
for some functions $\tilde{F}_{0}$, $\tilde{F}_{1}$ and $\tilde{G}$.
The corresponding MFG cost functional is given by,
$J_{g}(X_{0},\alpha,\mu)=\mathbb{E}\left[\int_{0}^{T}\left[\tilde{F}_{0}(t,X^{\alpha}_{t},\alpha_{t})+\delta_{\mu}\tilde{F}_{1}(t,\mu_{t})(X^{\alpha}_{t})\right]\,\mathrm{d}t+\delta_{\mu}\tilde{G}(\mu_{T})(X^{\alpha}_{T})\right],$
This is exactly the MFG studied in many of the earlier studies such as [4]. In particular, it is a standard MFG as the mean-field interaction is only through the law of the current state and not the control. Then, the measure flow takes values in ${\mathscr{P}}_{2}(\mathbb{R}^{n})$ instead of ${\mathscr{P}}_{2}({\mathscr{D}}_{n})$ or ${\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)$. Further, by the continuity of $t\mapsto X^{\alpha}_{t}$, the flow $t\mapsto\mu_{t}\in{\mathscr{P}}_{2}(\mathbb{R}^{n})$ is continuous, and the Corollary 4.1 holds with a continuous deterministic flow $\mu^{*}_{t}={\mathcal{L}}(X^{\alpha^{*}}_{t})$.
4.3 Deterministic and Stochastic Formulations
Although the formulation of [4] is seemingly different than that of ours, they are equivalent and their results are included in our framework. Indeed, control problems on the space of measures have been studied in varying settings. In particular, [25, 26] provide a comprehensive overview of different stochastic formulations of the problem. In the absence of common noise and with Markovian dynamics, these problems also admit deterministic formulations that directly control the evolution of laws via the Fokker-Planck equation [3, 4, 5, 16, 49, 51, 57, 58].
To further explain the connection between these stochastic and deterministic formulations we specialize to the case considered in [4]. Namely, we assume that $J$ is given as in (4.3), and consider the simple dynamics without jumps and common noise,
$\mathrm{d}X^{\alpha}_{t}=\alpha_{t}\,\mathrm{d}t+\sigma\mathrm{d}W_{t}.$
In this setting these two formulations are the followings.
  1. (1)
    Stochastic (strong) formulation. On a given probabilistic basis, the objective is to minimize $J(X_{0},\alpha)$ of (4.3) over controls $\alpha\in{\mathscr{A}}$, with above dynamics.
  2. (2)
    Deterministic formulation. In this formulation the state process is the law of the controlled diffusions $\mu^{\alpha}_{t}={\mathcal{L}}(X^{\alpha}_{t})$, and the controls are of feedback type. Then, the analogous control problem is to minimize
    $\int_{0}^{T}[\,\int_{\mathbb{R}^{n}}\tilde{F}_{0}(t,x,\alpha(t,x))\,\mu^{\alpha}_{t}(\mathrm{d}x)\ +\ \tilde{F}_{1}(t,\mu^{\alpha}_{t})]\,\mathrm{d}t+\tilde{G}(\mu^{\alpha}_{T}),$
    over all measurable feedback functions $\alpha:[0,T]\times\mathbb{R}^{n}\mapsto U$, where the state $\mu^{\alpha}\in{\mathscr{P}}(\mathbb{R}^{d})$ solves the Fokker-Planck equation:
    $\partial_{t}\mu^{\alpha}+\mathrm{div}(\alpha\mu^{\alpha})-\frac{\sigma^{2}}{2}\Delta\mu^{\alpha}=0.$
    The above is understood in a distributional sense with initial condition $\mu^{\alpha}_{0}={\mathcal{L}}(X_{0})$.
The equivalence between between these two formulations is proved using the superposition principle that relates a solution of a Fokker–Planck equation to a weak solution of a SDE, see for example [45] which establish this fact under much more general dynamics. Starting from the deterministic formulation, this allows us to define a weak control for the stochastic problem in feedback form with the probability space being part of the control. One then appeals to a mimicking theorem to establish the equivalence of weak control processes and weak controls in feedback form. Additional arguments such as boundedness of feedback controls are utilized to obtain a strong control on a fixed probability space. In the present setting, a detailed implementation of these steps are given in [16, Proposition 2.2].
5 Proof of Theorem 3.6
The main idea is to compose the optimal control $\alpha^{*}$ with an arbitrary control $\alpha$ through a biased independent coin-flip. To achieve this, we extend the original probabilistic structure by adding an initial Bernoulli random variable, and then we use either one of the controls depending on its outcome. This provides linearity in the distribution variables and we obtain the result by essentially differentiating with respect to success probability of the coin-flip. The assumed invariance principle ensures that $\alpha^{*}$ is still a minimizer of the extended problem.
5.1 Preliminaries
We first extend the probability basis $\gamma$ to $\hat{\gamma}$ as follows,
$\hat{\Omega}:=\Omega\times\{0,1\},\quad\hat{\mathscr{F}}:={\mathscr{F}}\otimes{\mathscr{F}}^{\diamond},\quad\hat{\mathbb{P}}:=\mathbb{P}\otimes\tfrac{1}{2}(\delta_{\{0\}}+\delta_{\{1\}}),\quad\hat{\mathscr{F}}_{t}:={\mathscr{F}}_{t}\otimes{\mathscr{F}}^{\diamond},$
where ${\mathscr{F}}^{\diamond}$ is the power set of $\{0,1\}$. We naturally extend any function $\zeta$ on $\Omega$ to a function on $\hat{\Omega}$ by setting $\hat{\zeta}(\omega,x):=\zeta(\omega)$, $(\omega,x)\in\hat{\Omega}$. Then, $\hat{\gamma}:=(\hat{\Omega},\hat{\mathscr{F}},\hat{\mathbb{P}},\hat{\mathbb{F}},\hat{\mathscr{W}},\hat{\mathscr{N}})$ is a probability basis. Additionally, the $\sigma$-algebra $\hat{\mathscr{G}}$ generated by the common noise $(\hat{W}^{0},\hat{N}^{0})$ is given by
$\hat{\mathscr{G}}={\mathscr{G}}\otimes{\mathscr{G}}^{\diamond},$
where ${\mathscr{G}}^{\diamond}:=\{\emptyset,\{0,1\}\}$ is the trivial $\sigma$-algebra. Given any two functions $\zeta,\eta$ on $\Omega$, we define a new function $\zeta\otimes\eta$ on $\hat{\Omega}$ by,
$(\zeta\otimes\eta)(\omega,x):=\zeta(\omega)\chi_{\{x=0\}}+\eta(\omega)\chi_{\{x=1\}},\qquad(\omega,x)\in\hat{\Omega}.$
Then, for any $Z=Z(\omega,x)$, $\hat{Z}=Z\otimes Z$ and $Z=Z^{(0)}\otimes Z^{(1)}$ where $Z^{(x)}(\omega)=Z(\omega,x)$.
Given $\delta\in(0,1)$, we define a probability measure $\mathbb{P}^{\delta}$ equivalent to $\hat{\mathbb{P}}$ by,
$\mathbb{E}^{\delta}[Z]=\mathbb{E}^{\mathbb{P}^{\delta}}[Z]:=\delta\,\mathbb{E}[Z^{(0)}]+(1-\delta)\,\mathbb{E}[Z^{(1)}].$ (5.1)
This defines a family of probability bases $\gamma^{\delta}=(\hat{\Omega},\hat{\mathscr{F}},\mathbb{P}^{\delta},\hat{\mathbb{F}},\hat{\mathscr{W}},\hat{\mathscr{N}})$ indexed by $\delta\in(0,1)$. As $\mathbb{P}^{\delta}$ is equivalent to $\hat{\mathbb{P}}$, we have $\hat{\mathscr{A}}:={\mathscr{A}}(\hat{\gamma})={\mathscr{A}}(\gamma^{\delta})$. For $\beta\in\hat{\mathscr{A}}$, the projected controls $\beta^{(0)},\beta^{(1)}$ are in ${\mathscr{A}}$ and by definition $\beta_{t}=\beta^{(0)}_{t}\otimes\beta^{(1)}_{t}$. Let $Y^{\beta}$ be the solution to $\mathrm{SDE}^{\hat{\gamma}}(\hat{X}_{0},\beta)$, and $X^{(0)}:=X^{\beta^{(0)}}$, $X^{(1)}:=X^{\beta^{(1)}}$ be the controlled state processes solving the equation SDE$(X_{0},\alpha)$ with controls $\beta^{(0)}$ and $\beta^{(1)}$, respectively. Further, we set,
$\nu^{\beta,\delta}_{t}:={\mathcal{L}}^{\mathbb{P}^{\delta}}((Y^{\beta}_{\cdot\land t},\beta_{t})\,|\,\hat{\mathscr{G}}\,),\qquad\nu^{(x)}_{t}:={\mathcal{L}}^{\mathbb{P}}((X^{(x)}_{\cdot\land t},\beta^{(x)}_{t})\,|\,{\mathscr{G}}\,),\qquad x=0,1.$
Since $\hat{\mathscr{G}}={\mathscr{G}}\otimes{\mathscr{G}}^{\diamond}$, $\nu^{\beta,\delta}$ is identified with a function on $\Omega$.
We have the following immediate consequences of the definitions.
Lemma 5.1.
For all $t\in[0,T]$, a.s., $Y_{t}^{\beta}=X_{t}^{(0)}\otimes X_{t}^{(1)}$ and $\nu^{\beta,\delta}_{t}=\delta\nu^{(0)}_{t}+(1-\delta)\nu^{(1)}_{t}$.
Proof.
First consider the Itô-integral,
$I_{t}(\beta)(\omega,x)$ $:=(\int_{0}^{t}\sigma(s,Y^{\beta}_{\cdot\land s},\beta_{s})\,\mathrm{d}\hat{W}_{s})(\omega,x)$
$=(\int_{0}^{t}\sigma(s,X^{(0)}_{\cdot\land s},\alpha^{(0)}_{s})\chi_{\{x=0\}}\,\mathrm{d}\hat{W}_{s})(\omega,x)+(\int_{0}^{t}\sigma(s,X^{(1)}_{\cdot\land s},\alpha^{(1)}_{s})\chi_{\{x=1\}}\,\mathrm{d}\hat{W}_{s})(\omega,x)$
$=(\int_{0}^{t}\sigma(s,X^{(0)}_{\cdot\land s},\alpha^{(0)}_{s})\,\mathrm{d}W_{s})(\omega)\chi_{\{x=0\}}+(\int_{0}^{t}\sigma(s,X^{(1)}_{\cdot\land s},\alpha^{(1)}_{s})\,\mathrm{d}W_{s})(\omega)\chi_{\{x=1\}}$
[…]
where the final equality follows by properties of the stochastic integral, see for example [41, Exercise 2.30,Chapter 3.2]. Hence, $I_{t}(\beta)=I_{t}(\alpha^{(0)})\otimes I_{t}(\alpha^{(1)})$ a.s. A similar computation of all other four integrals in $\mathrm{SDE}^{\hat{\gamma}}(\hat{X}_{0},\beta)$ implies that $X^{(0)}\otimes X^{(1)}$ is a solution of $\mathrm{SDE}^{\hat{\gamma}}(\hat{X}_{0},\beta)$. Then by strong uniqueness, we conclude that $Y^{\beta}=X^{(0)}\otimes X^{(1)}$.
Let $\xi:(\omega,x)\in\hat{\Omega}\mapsto\xi(\omega,x)=x$ be the canonical projection. For $t\in[0,T]$ and a Borel set $S\subset{\mathscr{D}}_{n}\times U$, the followings hold a.s.,
$\nu^{\beta,\delta}(S)$ $=\mathbb{P}^{\beta}((Y^{\beta}_{\cdot\land t},\beta_{t})\in S\,|\,\hat{\mathscr{G}})$
$=\mathbb{P}^{\delta}((Y^{\beta}_{\cdot\land t},\beta_{t})\in S\ \text{and}\ \xi=0\,|\,\hat{\mathscr{G}})+\mathbb{P}^{\delta}((Y^{\beta}_{\cdot\land t},\beta_{t})\in S\ \text{and}\ \xi=1\,|\,\hat{\mathscr{G}})$
$=\mathbb{P}^{\delta}((X^{(0)}_{\cdot\land t},\alpha^{(0)}_{t})\in S\ \text{and}\ \xi=0\,|\,\hat{\mathscr{G}})+\mathbb{P}^{\delta}((X^{(1)}_{\cdot\land t},\alpha^{(1)}_{t})\in S\ \text{and}\ \xi=1\,|\,\hat{\mathscr{G}}).$
Since the canonical map $\xi$ is independent of all random elements defined on the original probability space $\Omega$ and $\hat{\mathscr{G}}={\mathscr{G}}\otimes{\mathscr{G}}^{\diamond}$, a.s.,
$\nu^{\beta,\delta}(S)(\omega,x)$ $=\mathbb{P}^{\delta}(\xi=0)\ \mathbb{P}((X^{(0)}_{\cdot\land t},\alpha^{(0)}_{t})\in S\,|\,{\mathscr{G}})+\mathbb{P}^{\delta}(\xi=1)\ \mathbb{P}((X^{(1)}_{\cdot\land t},\alpha^{(1)}_{t})\in S\,|\,{\mathscr{G}})$
$=\delta\,\nu^{(0)}(S)(\omega)+(1-\delta)\,\nu^{(1)}(S)(\omega).$
For a given $Y_{0}\in{\mathscr{L}}^{2}(\hat{\mathscr{F}}_{0})$, $\delta\in(0,1)$, and a ${\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)$-valued process $\eta$ on $\Omega$, set
${\mathcal{J}}_{\delta}(Y_{0},\beta,\eta)$ $:=\mathbb{E}^{\delta}\left[\int_{0}^{T}F(t,Y^{\beta}_{\cdot\wedge t},\beta_{t},\hat{\eta}_{t})\ \mathrm{d}t+G(Y^{\beta}_{\cdot\wedge T},\hat{\eta}_{T})\right],$
${\mathcal{J}}(X_{0},\alpha,\eta)$ $:=\mathbb{E}\left[\int_{0}^{T}F(t,X^{\alpha}_{\cdot\wedge t},\alpha_{t},\eta_{t})\ \mathrm{d}t+G(X^{\alpha}_{\cdot\wedge T},\eta_{T})\right].$
Then, the MFC problem on $\gamma^{\delta}$ is to minimize $J_{\delta}(Y_{0},\beta):={\mathcal{J}}_{\delta}(Y_{0},\beta,\nu^{\beta,\delta})$ over all $\beta\in\hat{\mathscr{A}}$. Moreover, in view of (5.1),
$J_{\delta}(Y_{0},\beta)=\delta\ {\mathcal{J}}(Y_{0}^{(0)},\beta^{(0)},\nu^{\beta,\delta})+(1-\delta)\ {\mathcal{J}}(Y_{0}^{(1)},\beta^{(1)},\nu^{\beta,\delta}).\,$ (5.2)
Additionally, as ${\mathcal{L}}^{\mathbb{P}^{\delta}}(\hat{X}_{0})={\mathcal{L}}^{\mathbb{P}}(X_{0})$, the assumed invariance principle implies that
$\inf_{\beta\in\hat{\mathscr{A}}}J_{\delta}(\hat{X}_{0},\beta)=\inf_{\alpha\in{\mathscr{A}}}J(X_{0},\alpha).$ (5.3)
5.2 Completion of the Proof
Let $\alpha^{*}\in{\mathscr{A}}$ be a minimizer of MFC (3.1) starting from initial condition $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}_{0})$, and let $\alpha\in{\mathscr{A}}$ be an arbitrary control process. To simplify the presentation, we write $X^{*}=X^{\alpha^{*}}$ for the optimal state process and
$\nu^{\alpha}_{t}:={\mathcal{L}}^{\alpha}_{t}={\mathcal{L}}((X^{\alpha}_{\cdot\wedge t},\alpha_{t})\,|\,{\mathscr{G}}\,),\qquad\nu^{*}_{t}:={\mathcal{L}}^{\alpha^{*}}_{t}={\mathcal{L}}((X^{*}_{\cdot\wedge t},\alpha^{*}_{t})\,|\,{\mathscr{G}}\,).$
Set $\beta:=\alpha\otimes\alpha^{*}$, $Y_{0}=X_{0}\otimes X_{0}=\hat{X}_{0}$, so that $\beta^{(0)}=\alpha$, $\beta^{(1)}=\alpha^{*}$, $Y_{0}^{(0)}=Y_{0}^{(1)}=X_{0}$, and by Lemma 5.1, $\nu^{\beta,\delta}_{t}=\delta\nu^{\alpha}_{t}+(1-\delta)\nu^{*}_{t}$. Moreover, by (5.3), $J(X_{0},\alpha^{*})\leq J_{\delta}(Y_{0},\beta)=J_{\delta}(\hat{X}_{0},\beta)$.
The following calculation is central to the proof.
Lemma 5.2.
Let $J_{g}$ be as in (3.3). Then,
$0\leq\lim_{\delta\downarrow 0}\frac{1}{\delta}\left(J_{\delta}(\hat{X}_{0},\beta)-J(X_{0},\alpha^{*})\right)=J_{g}(\alpha,X_{0},\nu^{*})-J_{g}(X_{0},\alpha^{*},\nu^{*}).$
In particular, $\alpha^{*}$ minimizes $J_{g}(\alpha,X_{0},\nu^{*})$ over all $\alpha\in{\mathscr{A}}$.
Proof.
We first note that the growth conditions in Assumption 3.5 and the second moment estimate of the controlled state process in Theorem 2.4 justify the use of Fubini’s theorem in the rest of this proof.
By the definitions of $J_{\delta},{\mathcal{J}}_{\delta}$, and ${\mathcal{J}}$,
$J_{\delta}(\hat{X}_{0},\beta)$ $={\mathcal{J}}_{\delta}(\hat{X}_{0},\beta,\nu^{\beta,\delta})=\delta{\mathcal{J}}(X_{0},\alpha,\nu^{\beta,\delta})+(1-\delta){\mathcal{J}}(X_{0},\alpha^{*},\nu^{\beta,\delta}).$
Then, by (5.2),
$\frac{1}{\delta}\left(J_{\delta}(\hat{X}_{0},\beta)-J(X_{0},\alpha^{*})\right)={\mathcal{I}}_{1,\delta}+{\mathcal{I}}_{2,\delta},\quad\text{where}$
${\mathcal{I}}_{1,\delta}:={\mathcal{J}}(X_{0},\alpha,\nu^{\beta,\delta})-{\mathcal{J}}(X_{0},\alpha^{*},\nu^{\beta,\delta}),\quad{\mathcal{I}}_{2,\delta}:=\frac{1}{\delta}\left({\mathcal{J}}(X_{0},\alpha^{*},\nu^{\beta,\delta})-J(X_{0},\alpha^{*})\right).$
Since $\nu^{\beta,\delta}$ converges to $\nu^{*}$ as $\delta$ tends to zero,
$\lim_{\delta\downarrow 0}\ {\mathcal{I}}_{1,\delta}$ $={\mathcal{J}}(X_{0},\alpha,\nu^{*})-{\mathcal{J}}(X_{0},\alpha^{*},\nu^{*})$ (5.4)
$=\mathbb{E}\left[\int_{0}^{T}\left[F(t,X^{\alpha}_{\cdot\wedge t},\alpha_{t},\nu^{*}_{t})-F(t,X^{*}_{\cdot\wedge t},\alpha^{*}_{t},\nu^{*}_{t})\right]\mathrm{d}t+G(X^{\alpha}_{\cdot\wedge T},\nu^{*}_{T})-G(X^{*}_{\cdot\wedge T},\nu^{*}_{T})\right].$
To analyze the limit of ${\mathcal{I}}_{2,\delta}$ we first observe that by Lemma 5.1, a.s.,
$\lim_{\delta\downarrow 0}\frac{1}{\delta}[F(t,$ $X^{*}_{\cdot\wedge t},\alpha^{*}_{t},\nu^{\beta,\delta}_{t})-F(t,X^{*}_{\cdot\wedge t},\alpha^{*}_{t},\nu^{*}_{t})]$
$=$ $\lim_{\delta\downarrow 0}\ \int_{0}^{1}\int_{{\mathscr{D}}_{n}\times U}\delta_{\nu}F(t,X^{*}_{\cdot\wedge t},\alpha^{*}_{t},\tau\delta\nu^{\alpha}_{t}+(1-\tau\delta)\nu^{*}_{t})(x,a)\,(\nu^{\alpha}_{t}-\nu^{*}_{t})(\mathrm{d}x,\mathrm{d}a)\ \mathrm{d}\tau$
$=$ $\int_{{\mathscr{D}}_{n}\times U}\delta_{\nu}F(t,X^{*}_{\cdot\wedge t},\alpha^{*}_{t},\nu^{*}_{t})(x,a)\,(\nu^{\alpha}_{t}-\nu^{*}_{t})(\mathrm{d}x,\mathrm{d}a).$
We use the definition $\nu^{*}_{t}={\mathcal{L}}((X^{*}_{\cdot\wedge t},\alpha^{*}_{t})\,|\,{\mathscr{G}})$ and well-known properties of the conditional expectations, to compute:
$\mathbb{E}\left[\int_{{\mathscr{D}}_{n}\times U}\delta_{\nu}F(t,X^{*}_{\cdot\land t},\alpha^{*}_{t},\nu^{*}_{t})(x,a)\,\nu^{*}_{t}(\mathrm{d}x,\mathrm{d}a)\right]$
$\hskip 50.0pt=\mathbb{E}\left[\mathbb{E}\left[\int_{{\mathscr{D}}_{n}\times U}\delta_{\nu}F(t,X^{*}_{\cdot\land t},\alpha^{*}_{t},\nu^{*}_{t})(x,a)\,\nu^{*}_{t}(\mathrm{d}x,\mathrm{d}a)\,|\,{\mathscr{G}}\,\right]\right]$
$\hskip 50.0pt=\mathbb{E}\left[\int_{{\mathscr{D}}_{n}\times U}\int_{{\mathscr{D}}_{n}\times U}\delta_{\nu}F(t,\tilde{x},\tilde{a},\nu^{*}_{t})(x,a)\,\nu^{*}_{t}(\mathrm{d}x,\mathrm{d}a)\,\nu^{*}_{t}(\mathrm{d}\tilde{x},\mathrm{d}\tilde{a})\right]$
$\hskip 50.0pt=\mathbb{E}\left[\int_{{\mathscr{D}}_{n}\times U}\int_{{\mathscr{D}}_{n}\times U}\delta_{\nu}F(t,\tilde{x},\tilde{a},\nu^{*}_{t})(x,a)\,\nu^{*}_{t}(\mathrm{d}\tilde{x},\mathrm{d}\tilde{a})\,\nu^{*}_{t}(\mathrm{d}x,\mathrm{d}a)\right]$
$\hskip 50.0pt=\mathbb{E}\left[\mathbb{E}\left[\int_{{\mathscr{D}}_{n}\times U}\delta_{\nu}F(t,\tilde{x},\tilde{a},\nu^{*}_{t})(X^{*}_{\cdot\land t},\alpha^{*}_{t})\,\nu^{*}_{t}(\mathrm{d}\tilde{x},\mathrm{d}\tilde{a})\,\Bigl{|}\,{\mathscr{G}}\,\right]\right]$
$\hskip 50.0pt=\mathbb{E}\left[\int_{{\mathscr{D}}_{n}\times U}\delta_{\nu}F(t,\tilde{x},\tilde{a},\nu^{*}_{t})(X^{*}_{\cdot\land t},\alpha^{*}_{t})\,\nu^{*}_{t}(\mathrm{d}\tilde{x},\mathrm{d}\tilde{a})\right].$
As $\nu^{\alpha}_{t}={\mathcal{L}}((X^{\alpha}_{\cdot\wedge t},\alpha_{t})\,|\,{\mathscr{G}})$, a computation as above implies that
$\mathbb{E}\left[\int_{{\mathscr{D}}_{n}\times U}\delta_{\nu}F(t,X^{*}_{t},\alpha^{*}_{t},\nu^{*}_{t})(x,a)\,\nu^{\alpha}_{t}(\mathrm{d}x,\mathrm{d}a)\right]$
$\hskip 50.0pt=\mathbb{E}\left[\int_{{\mathscr{D}}_{n}\times U}\delta_{\nu}F(t,\tilde{x},\tilde{a},\nu^{*}_{t})(X^{\alpha}_{\cdot\land t},\alpha_{t})\,\nu^{*}_{t}(\mathrm{d}\tilde{x},\mathrm{d}\tilde{a})\right].$
Above calculations, (5.4) and the definition of $f$ in (1.4) imply that
$\lim_{\delta\downarrow 0}\ {\mathcal{I}}_{1,\delta}+\frac{1}{\delta}$ $\mathbb{E}\int_{0}^{T}\left[F(t,X^{*}_{\cdot\wedge t},\alpha^{*}_{t},\nu^{\beta,\delta}_{t})-F(t,X^{*}_{\cdot\wedge t},\alpha^{*}_{t},\nu^{*}_{t})\right]\mathrm{d}t$
$\hskip 30.0pt=\mathbb{E}\int_{0}^{T}\left[f(t,X^{\alpha}_{\cdot\wedge t},\alpha_{t},\nu^{*}_{t})-f(t,X^{*}_{\cdot\wedge t},\alpha^{*}_{t},\nu^{*}_{t})\right]\mathrm{d}t.$
Similarly, with $g$ as in (1.4),
$\lim_{\delta\downarrow 0}\ \mathbb{E}\left[G(X^{\alpha}_{\cdot\wedge T},\nu^{\beta,\delta}_{T})-G(X^{*}_{\cdot\wedge T},\nu^{\beta,\delta}_{T})\right]=\mathbb{E}\left[g(X^{\alpha}_{\cdot\wedge T},\nu^{*}_{T})-g(X^{*}_{\cdot\wedge T},\nu^{*}_{T})\right].$
We have shown that $\alpha^{*}$ minimizes $J_{g}(X_{0},\alpha,\nu^{*})$ and by its definition $\nu^{*}_{t}={\mathcal{L}}((X^{*}_{\wedge t},\alpha^{*}_{t})\,|\,{\mathscr{G}})$. Hence, $(\alpha^{*},\nu^{*})$ is a Nash equilibrium of MFG of controls.
$\Box$
Remark 5.3.
We use the invariance property only in (5.3). Therefore it suffices to require that value functions on $\gamma$ and its extension $\hat{\gamma}$ have the same value functions. In fact, if the probability basis $\gamma$ is rich enough in the sense defined in [20, Section 2.1.3] and $X_{0}$ is deterministic, one can directly prove that the extension to $\hat{\gamma}$ does not change the value.
Additionally, if the triplet $(\gamma,X_{0},\alpha^{*})$ minimizes the ‘global’ MFC problem (3.2), then above proof implies that $(\alpha^{*},\nu^{*})$ is a Nash equilibrium on the basis $\gamma$ starting from $X_{0}$.
6 Examples and Connection to Gradient Flows
In this section we propose a general approach that connects the Langevin equation (1.1) to MFC and MFG and illustrate this approach with three examples.
6.1 General Approach
Consider the solution of the McKean-Vlasov stochastic differential equations of the form (1.1), with a given functional $F:{\mathscr{P}}_{2}(\mathbb{R}^{n})\to\mathbb{R}$ on some filtered probability space $(\Omega,{\mathscr{F}},\mathbb{F},\mathbb{P})$ supporting a Brownian motion $W$ and a fixed volatility $\sigma>0$. As discussed in the Introduction, equations of the type (1.1) are called mean-field Langevin dynamics, and the corresponding non-linear Fokker-Planck equation can be interpreted as a gradient flow of measures with respect to the $2$-Wasserstein distance.
For a linearly convex functional $F$ , [34] proves that the laws of the solution to (1.1) converge to the minimizer of the (strictly convex) free energy
$F(\mu)+\frac{\sigma^{2}}{2}H(\mu),$
where $H(\mu)$ is the relative entropy of $\mu\in{\mathscr{P}}_{2}(\mathbb{R}^{n})$ with respect to Lebesgue measure.
In the context of mean-field control or games, an alternative connection to an optimization problem is proposed for several models, including [14, 60, 61] for the Kuramoto synchronization and [54, 55] for the Cucker-Smale flocking. In general, we propose the following MFC with a separable structure,
$v(\mu):=\,\inf_{\alpha}\quad\mathbb{E}\left[\int_{0}^{\infty}e^{-\beta t}\left(\frac{1}{2}|\alpha_{t}|^{2}+F({\mathcal{L}}(X^{\alpha}_{t}))\right)\,\mathrm{d}t\right],\quad\mu\in{\mathscr{P}}_{2}(\mathbb{R}^{n}),$ (6.1)
where $dX^{\alpha}_{t}=\alpha_{t}\,\mathrm{d}t+\sigma\,\mathrm{d}W_{t}$, $\mu={\mathcal{L}}(X_{0})$ is given, and $\beta>0$ is a positive discount factor. It is standard that $v$ solves the following dynamic programming equation
$\beta v(\mu)=-\frac{1}{2}\mu\left(|\nabla_{x}\delta_{\mu}v(\mu)(\cdot)|^{2}\right)+\frac{1}{2}\mu\left(\Delta_{x}\delta_{\mu}v(\mu)(\cdot)\right)+F(\mu).$
Whenever the value function is smooth, the optimally controlled state process follows
$\mathrm{d}X_{t}=-\nabla_{x}\delta_{\mu}v({\mathcal{L}}(X_{t}))(X_{t})\,\mathrm{d}t+\sigma\,\mathrm{d}W_{t}.$ (6.2)
This is again an equation of type (1.1), except the functional $F$ is replaced by the value function $v$. Now, a standard control argument shows that linear convexity of $F$ implies linear convexity of the value function $v$, suggesting a similar structure of (1.1) and (6.2). Further, note that if $F$ is sufficiently smooth, one can always define a suitable control problem such that the optimally controlled state dynamics exactly match (1.1). Indeed, this can be achieved by replacing $F$ by,
$\widetilde{F}(\mu):=\beta F(\mu)+\frac{1}{2}\mu\left(|\nabla_{x}\delta_{\mu}F(\mu)(\cdot)|^{2}\right)-\frac{1}{2}\mu\left(\delta_{\mu}F(\mu)(\cdot)\right).$
Then, a direct calculation shows that the value function with running cost $\widetilde{F}$ is exactly $F$, and the optimal state process solves the mean-field Langevin equation (1.1).
Using our main result, minimizers of the control problem give rise to Nash equilibria of an associated MFG with cost functional,
$J_{g,\beta}(X_{0},\alpha,\mu^{*}):=\mathbb{E}\left[\int_{0}^{\infty}e^{-\beta t}\left(\frac{1}{2}|\alpha_{t}|^{2}+\delta_{\mu}F(\mu^{*}_{t})(X^{\alpha}_{t})\right)\,\mathrm{d}t\right],$ (6.3)
and state dynamics $X^{\alpha}_{t}=X_{0}+\int_{0}^{t}\alpha_{s}\,\mathrm{d}s+\sigma W_{t}$. We continue to illustrating the above discussion with two prominent examples of mean-field games.
6.2 Kuramoto Synchronization
The classical Kuramoto dynamical system is a fascinating phenomenological model for collective synchronization [43]. It postulates the following system of coupled ordinary differential equations describing the phases $\theta$ of $N$ oscillators,
$\frac{\mathrm{d}}{\mathrm{d}t}\theta^{k}_{t}=\omega^{k}+\frac{\kappa}{N}\sum_{j=1}^{N}\sin(\theta^{j}_{t}-\theta^{k}_{t}),\quad k=1,..,N.$
Here $\omega^{k}$ is the natural frequency of the $k$-th oscillator, and $\kappa>0$ denotes the coupling strength, or the interaction parameter. The survey article [59] provides an excellent overview and account of the history of this rich subject.
An intruging property of this model is a phase transition across a critical value of $\kappa$, from incoherence to spontaneous synchronization. The related mean-field model studied in [28] that shares this feature is a McKean-Vlasov equation obtained by setting $\omega^{k}$ to zero and adding a Brownian noise with strength $\sigma>0$. The resulting dynamics read,
$\mathrm{d}X_{t}=-\kappa\int_{\mathbb{T}}\sin(X_{t}-y)\,{\mathcal{L}}(X_{t})(\mathrm{d}y)\,\mathrm{d}t+\sigma\,\mathrm{d}W_{t},$
with the one-dimensional torus $\mathbb{T}$ as its state space. This flow is in the form of a mean-field Langevin SDE (1.1), for the potential
$F(\mu):=\int_{\mathbb{T}}\int_{\mathbb{T}}\sin^{2}((x-y)/2)\,\mu(\mathrm{d}x)\,\mu(\mathrm{d}y)\quad\Rightarrow\quad\delta_{\mu}F(\mu,x)=\int_{\mathbb{T}}2\sin^{2}((x-y)/2)\mu(\mathrm{d}y).$
Then, using the trigonometric identity $2\sin^{2}(x/2)=1-\cos(x)$, we compute that
$\nabla_{x}\delta_{\mu}F(\mu,x)=\int_{\mathbb{T}}\nabla_{x}(1-\cos(x-y))\mu(\mathrm{d}y)=\int_{\mathbb{T}}\sin(x-y)\mu(\mathrm{d}y).$
The associated MFC problem minimizes
$\inf_{\alpha}\quad\mathbb{E}\int_{0}^{\infty}e^{-\beta t}\left(\frac{1}{2}\alpha_{t}^{2}+F({\mathcal{L}}(X^{\alpha}_{t}))\right)\,\mathrm{d}t$
subject to $X^{\alpha}_{t}=\int_{0}^{t}\alpha_{s}\,\mathrm{d}s+\sigma W_{t}$. In view of Theorem 3.6, any minimizer $\alpha^{*}$ of this MFC gives rise to a mean-field Nash equilibrium $\mu^{*}_{t}={\mathcal{L}}(X^{\alpha^{*}}_{t})$, in which players minimize,
$\inf_{\alpha}\quad\mathbb{E}\int_{0}^{\infty}e^{-\beta t}\left(\frac{1}{2}\alpha_{t}^{2}+\kappa\int_{\mathbb{T}}2\sin^{2}\left(\frac{X^{\alpha}_{t}-y}{2}\right)\,\mu^{*}_{t}(\mathrm{d}y)\right)\,\mathrm{d}t.$
The above MFG approach to synchronization was first proposed by Yin, Mehta, Meyn, & Shanbhag in [61, 60], and later used by Carmona & Graves [12] to study jet-lag recovery by modeling the alignment with the circadian rhythm. These studies and the recent works of Carmona, Cormier, & Soner [14], Cesaroni & Cirant [18], and Höfer & Soner [33] establish phase transitions in the Kuramoto MFGs analogous to the one exhibited by the original and mean-field Kuramoto dynamical system, providing evidence for a connection between these seemingly very different models.
6.3 Cucker-Smale Flocking
In their classical paper [21], Cucker & Smale propose a model of a flock of $N$ birds in which the state of the $k$-th bird is described by a vector $(x^{k},v^{k})\in\mathbb{R}^{3}\times\mathbb{R}^{3}$ where $x^{k}$ denotes the location and $v^{k}$ the velocity. With given constants $\kappa>0$, $\rho\geq 0$, they postulate,
$\frac{\mathrm{d}}{\mathrm{d}t}x^{k}_{t}=v^{k}_{t},\qquad\frac{\mathrm{d}}{\mathrm{d}t}v^{k}_{t}=\frac{\kappa}{N}\sum_{k=1}^{N}\frac{(v^{j}_{t}-v^{k}_{t})}{\phi(x_{t}-\tilde{x}_{t})},\qquad\phi(x):=(1+|x|^{2})^{\rho}.$ (6.4)
Flocking is interpreted as $\max_{k,j=1,..,N}|x^{k}_{t}-x^{j}_{t}|$ staying bounded for all times and [21] proves that flocking occurs if $\rho<1/2$. Motivated by this model, Nourian, Caines & Malhamé [54, 55] propose an ergodic mean-field game in which the representative player’s running cost is given by,
$f((x,v),a,\mu)=\frac{1}{2}|a|^{2}+\frac{\kappa}{2}\Bigl{|}\int_{\mathbb{R}^{2n}}\frac{v-\tilde{v}}{\phi(x-\tilde{x})}\,\mu(\mathrm{d}\tilde{x},\mathrm{d}\tilde{v})\Bigr{|}^{2},$
where $a\in\mathbb{R}^{n}$, $\mu\in{\mathscr{P}}_{2}(\mathbb{R}^{2n})$. Also, [55] observes that $\rho=0$ reduces to a classical linear-quadratic control problem and performs a perturbation analysis for small $\rho$.
Although this system is not always a gradient flow, the velocity equation has the form
$\frac{\mathrm{d}}{\mathrm{d}t}v^{k}_{t}=-\ \frac{\partial}{\partial v^{k}}\delta_{\mu}F(\mu^{N}_{t})(x^{k}_{t},v^{k}_{t}),$ (6.5)
where $\mu^{N}_{t}$ is the empirical measure of $((x^{1}_{t},v^{1}_{t}),..,(x^{N}_{t},v^{N}_{t}))$ and
$F(\mu):=\frac{\kappa}{4}\int_{\mathbb{R}^{2n}}\int_{\mathbb{R}^{2n}}\frac{|v-\tilde{v}|^{2}}{\phi(x-\tilde{x})}\,\mu(\mathrm{d}x,\mathrm{d}v)\,\mu(\mathrm{d}\tilde{x},\mathrm{d}\tilde{v}),\quad\mu\in{\mathscr{P}}_{2}(\mathbb{R}^{2n}).$
Indeed, we calculate that
$\delta_{\mu}F(\mu)(x,v)=\frac{\kappa}{2}\int_{\mathbb{R}^{2n}}\frac{|v-\tilde{v}|^{2}}{\phi(x-\tilde{x})}\,\mu(\mathrm{d}\tilde{x},\mathrm{d}\tilde{v})\ \Rightarrow\ \nabla_{v}\delta_{\mu}F(\mu)(x,v)=\kappa\int_{\mathbb{R}^{2n}}\frac{v-\tilde{v}}{\phi(x-\tilde{x})}\,\mu(\mathrm{d}\tilde{x},\mathrm{d}\tilde{v}).$
Consistent this observation, [11, Chapter 4.7.3] suggests a slightly different MFG than that of [54]. It is a potential game corresponding to the MFC of (6.1) with running cost $F$ and the $2n$-dimensional state dynamics
$\mathrm{d}X^{\alpha}_{t}=V^{\alpha}_{t}\,\mathrm{d}t,\quad\mathrm{d}V^{\alpha}_{t}=\alpha_{t}\,\mathrm{d}t+\sigma\,\mathrm{d}W_{t}.$
As in subsection 4.1, the corresponding MFG minimizes (6.3) with the running cost,
$f((x,v),a,\mu)=\frac{1}{2}|a|^{2}+\nabla_{v}\delta_{\mu}F(\mu)(x,v)=\frac{1}{2}|a|^{2}+\kappa\int_{\mathbb{R}^{2n}}\frac{v-\tilde{v}}{\phi(x-\tilde{x})}\,\mu(\mathrm{d}\tilde{x},\mathrm{d}\tilde{v}).$
In particular, if the value function $v(\mu)$ of this problem is smooth, then the optimal feedback control is given by $-\nabla_{v}\delta_{\mu}v(\mu)(x,v)$. Thus, the optimally controlled state dynamics follow
$dX^{*}_{t}=V^{*}_{t}\,\mathrm{d}t,\qquad dV^{*}_{t}=-\nabla_{v}\delta_{\mu}v(\mu_{t}^{*})(X^{*}_{t},V^{*}_{t})\,\mathrm{d}t+\sigma\,\mathrm{d}W_{t}.$
As in the pure gradient flow cases, above equation resembles closely the original model (6.5) with the value function replacing the functional $F$. Additionally, Theorem 3.6 implies that the law of $(X^{*},V^{*})$ is also a Nash equilibrium of the potential MFG (6.3).
6.4 Linear Quadratic Problem
The special case $\rho=0$ of the Cucker-Smale model is the well-known linear quadratic problem, which we now discuss. In this case, the mean-field limit potential is given by,
$F(\mu)=\frac{1}{2}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}|x-y|^{2}\,\mu(\mathrm{d}x)\,\mu(\mathrm{d}y)=\frac{1}{2}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}(|x|^{2}+|y|^{2}-2x.y)\,\mu(\mathrm{d}x)\mu(\mathrm{d}y)=\mathrm{Var}(\mu),$
for $\mu\in{\mathscr{P}}_{2}(\mathbb{R}^{n})$. Then, with $\mathfrak{m}(\mu):=\int y\,\mu(\mathrm{d}y)$,
$\delta_{\mu}F(\mu)(x)=\int_{\mathbb{R}^{n}}|x-y|^{2}\mu(\mathrm{d}y)\quad\Rightarrow\quad\nabla_{x}\delta_{\mu}F(\mu)(x)=2\int_{\mathbb{R}^{n}}(x-y)\mu(\mathrm{d}y)=2(x-\mathfrak{m}(\mu)).$
This implies that the corresponding mean-field Langevin equation is given by,
$\mathrm{d}X_{t}=-2(X_{t}-\mathbb{E}[X_{t}])\,\mathrm{d}t+\sigma\,\mathrm{d}W_{t}.$
The associated MFC with interaction parameter $\kappa>0$ is
$v(\mu):=\inf_{\alpha\in\mathcal{A}}\quad\mathbb{E}\int_{0}^{\infty}e^{-\beta t}\left(\frac{1}{2}|\alpha_{t}|^{2}+\kappa\,F(\mathcal{L}(X^{\alpha}_{t}))\right)\,\mathrm{d}t,$
where $X^{\alpha}_{t}=X_{0}+\int_{0}^{t}\alpha_{s}\,\mathrm{d}s+\sigma W_{t}$, and $\mu={\mathcal{L}}(X_{0})$. This is a simple linear, quadratic control problem and can be solved explicitly. Indeed, the dynamic programming equation is,
$\beta v(\mu)=-\frac{1}{2}\mu\left(|\nabla_{x}\delta_{\mu}v(\mu)(\cdot)|^{2}\right)+\frac{\sigma^{2}}{2}\mu\left(\Delta_{x}\delta_{\mu}v(\mu)(\cdot)\right)+\kappa\mathrm{Var}(\mu).$
The unique solution is $v(\mu)=a\mathrm{Var}(\mu)+b$ with,
$a=\frac{\sqrt{8\kappa+\beta^{2}}-\beta}{4},\quad b=\frac{\sigma^{2}na}{\beta},$
so that the optimally controlled state process $X^{*}$ follows,
$\mathrm{d}X^{*}_{t}=-2a(X^{*}_{t}-\mathbb{E}[X^{*}_{t}])\,\mathrm{d}t+\sigma\,\mathrm{d}W_{t}.$
The stationary measures of this equation are Gaussian distributions with arbitrary mean and variance $\sigma^{2}/(4a)$. The associated potential mean-field game minimizes
$J_{g,\beta}(X_{0},\alpha,\mu)=\mathbb{E}\int_{0}^{\infty}e^{-\beta t}\left(\frac{1}{2}|\alpha_{t}|^{2}+\kappa\int_{\mathbb{R}}|X^{\alpha}_{t}-y|^{2}\,\mu_{t}(\mathrm{d}y)\right)\,\mathrm{d}t.$
for a given continuous measure flow $t\in[0,\infty)\mapsto\mu_{t}\in{\mathscr{P}}_{2}(\mathbb{R}^{n})$. This mean-field game can be solved explicitly, and a direct computation shows that any Gaussian distribution with variance $\sigma^{2}/(4a)$ provides a stationary mean-field game equilibrium.
7 Invariance Principle
In this section, we prove the invariance principle stated in Definition 3.2 under under the following regularity and growth conditions. Let $\mathfrak{m}_{2}$ be the 2-Wasserstein metric on ${\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)$ and set ${\mathscr{X}}:={\mathscr{D}}_{n}\times U\times{\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)$. Recall the sup-norm $\|\cdot\|_{T}$ on ${\mathscr{D}}_{n}$, the notation ${\mathfrak{s}}(\nu)$ for the second moment (3.5), and the value function $v(X_{0},\gamma)$ of (3.1).
Assumption 7.1.
In addition Assumption 2.3, the followings hold:
  • (Local Uniform Boundedness). For any $\ell>0$, there exists a constant $c_{\ell}$ such that for any $(t,x,\zeta)\in[0,T]\times{\mathscr{D}}_{n}\times E$ and $|a|,|a^{\prime}|\leq\ell$,
    $|(b,\sigma,\sigma^{0})(t,x,a)-(b,\sigma,\sigma^{0})(t,x,a^{\prime})|$ $\leq c_{\ell},$
    $|(\lambda,\lambda^{0})(t,x,a,\zeta)-(\lambda,\lambda^{0})(t,x,a^{\prime},\zeta)|$ $\leq c_{\ell}L(\zeta).$
  • (Quadratic Growth). $F,G$ are measurable and there is a constant $c_{*}$ such that
    $|F(t,x,a,\nu)|+|G(x,\nu)|\leq C(1+\|x\|_{T}^{2}+|a|^{2}+\mathfrak{s}(\nu)),$
    for every $(t,x,a,\nu)\in[0,T]\times{\mathscr{X}}$.
  • (Uniform Continuity). For any $K>0$, there is a continuous map $m_{K}:[0,\infty)\mapsto[0,4K]$ with $m_{K}(0)=0$ such that for all $(t,x,\hat{x},\nu,\hat{\nu},a,\hat{a})\in[0,T]\times{\mathscr{X}}^{2}$,
    $|F_{K}(t,x,a,\nu)-F_{K}(t,\hat{x},\hat{a},\hat{\nu})|+|G_{K}(x,\nu)-G_{K}(\hat{x},\hat{\nu})|\leq m_{K}(\|x-\hat{x}\|_{T}+\mathfrak{m}_{2}(\nu,\hat{\nu})+|a-\hat{a}|),$
    where $F_{K}:=(F\wedge K)\vee(-K)$, $G_{K}:=(G\wedge K)\vee(-K)$.
  • (Feasibility). $v(X_{0},\gamma)<\infty$ for every $\gamma$ and $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}_{0}^{\gamma})$.
Theorem 7.2.
If the coefficients of an MFC problem satisfy Assumptions 2.3 and 7.1, then it is invariant under probability bases.
Remark 7.3.
In the diffusion setting, the invariance is proved in [20, Theorem 3.6] for ‘rich enough’ probabilistic structures [20, Section 2.1.3] and in [26, Proposition 2.4]. The counter example provided in [20, Section 3] shows the necessity of continuity assumptions. A careful analysis of the proof of our results shows that local uniform boundedness assumption can be replaced by ${\mathscr{L}}^{p}$ integrability of $X_{0}$ for some $p>2$, a condition used in several places, for example [44], to overcome similar difficulties.
This section is devoted to the proof of the above Theorem. Let ${\mathscr{C}}_{m}$ be the set of $\mathbb{R}^{m}$ valued continuous functions on $[0,T]$, and recall the measurable spaces $(E,{\mathscr{E}})$ and $(M,{\mathscr{M}})$ related to the point processes given in Appendix A.
For $\varrho\in{\mathscr{P}}_{2}(\mathbb{R}^{n})$ consider the canonical probability basis
$\gamma^{c}(\varrho)=(\Omega^{c},{\mathscr{F}}^{c},\mathbb{P}^{c},\mathbb{F}^{c},{\mathscr{W}}^{c},{\mathscr{N}}^{c}),$
where
  • $\Omega^{c}=\mathbb{R}^{n}\times{\mathscr{C}}_{\ell}\times{\mathscr{C}}_{d}\times M^{2}$;
  • $\mathbb{P}^{c}$ is probability measure on $\Omega^{c}$ given as the product of $\varrho$, the law of a standard $(\ell+d)$-dimensional Brownian motion, and the law of the point process on $M^{2}$;
  • The $\sigma$-algebra ${\mathscr{F}}^{c}$ is the $\mathbb{P}^{c}$ completion of ${\mathscr{B}}(\mathbb{R}^{n})\otimes{\mathscr{B}}({\mathscr{C}}_{\ell+d})\otimes{\mathscr{M}}^{\otimes 2}$;
  • ${\mathscr{W}}^{c}(\omega^{c})=(w^{0},w)$, ${\mathscr{N}}^{c}(\omega^{c})=(\pi^{0},\pi)$ for $\omega^{c}=(x_{0},w^{0},w,\pi^{0},\pi)\in\Omega^{c}$;
  • The filtration $\mathbb{F}^{c}=({\mathscr{F}}^{c}_{t})_{t\geq 0}$ is the right-continuous $\mathbb{P}^{c}$-completion of $\mathbb{F}^{\circ}$, where ${\mathscr{F}}^{\circ}_{t}$ is the smallest $\sigma$-algebra such that
    $\omega^{c}=(x_{0},w^{0},w,\pi^{0},\pi)\in\Omega^{c}\mapsto(x_{0},w^{0}_{\cdot\land t},w_{\cdot\land t},\pi^{0}(B),\pi(B))\in\mathbb{R}^{n+\ell+d+2k}$
    is measurable for every $B=[0,s]\times S$ where $s\leq t$, $S\in{\mathscr{E}}$.
We further set $X_{0}^{c}(\omega^{c}):=x_{0}$ for $\omega^{c}=(x_{0},w^{0},w,\pi^{0},\pi)\in\Omega^{c}$ so that ${\mathcal{L}}^{\mathbb{P}^{c}}(X_{0}^{c})=\varrho$.
For a basis $\gamma$, $t\in[0,T]$, and $\epsilon>0$, the following definitions are used in the analysis,
$\Xi^{\gamma}_{t}:=({\mathscr{W}}^{\gamma}_{s\wedge t},{\mathscr{N}}^{\gamma}_{s\wedge t})_{s\in[0,T]},\quad\Xi^{\gamma,\epsilon}_{t}:=(W^{\gamma}_{{s\vee\epsilon}\wedge t}-W^{\gamma}_{\epsilon},W^{0,\gamma}_{s\wedge t},{\mathscr{N}}^{\gamma}_{s\wedge t})_{s\in[0,T]},$ (7.1)
where ${\mathscr{N}}^{\gamma}_{\cdot\wedge t}$ refers to the restriction of ${\mathscr{N}}^{\gamma}$ to $[0,t]\times E$.
The following is an immediate consequence of the definition.
Lemma 7.4.
For any $\varrho\in{\mathscr{P}}_{2}(\mathbb{R}^{n})$, $\gamma\in\Gamma(\rho)$, and $X_{0}\in{\mathcal{I}}^{\gamma}(\varrho)$, $v(X_{0},\gamma)\leq v(X_{0}^{c},\gamma^{c}(\varrho))$.
Proof.
For $\alpha^{c}\in{\mathscr{A}}(\gamma^{c}(\varrho))$, we define $\alpha^{\gamma}_{t}(\omega):=\alpha^{c}(X_{0}(\omega),{\mathscr{W}}^{\gamma}(\omega),{\mathscr{N}}^{\gamma}(\omega))$ for $t\in[0,T]$, $\omega\in\Omega^{\gamma}$. It is clear that $\alpha\in{\mathscr{A}}(\gamma)$ and ${\mathcal{L}}^{\mathbb{P}^{\gamma}}(X_{0},{\mathscr{W}}^{\gamma},{\mathscr{N}}^{\gamma},\alpha^{\gamma})={\mathcal{L}}^{\mathbb{P}^{c}}(X_{0}^{c},{\mathscr{W}}^{c},{\mathscr{N}}^{c},\alpha^{c})$. Hence, by their definitions $J^{\gamma}(X_{0},\alpha^{\gamma})=J^{\gamma^{c}(\varrho)}(X_{0}^{c},\alpha^{c})$. Since for each control in the canonical setting we have constructed a control in $\gamma$ with same cost, this implies the claimed inequality. ∎
In order to prove Theorem 7.2, in the next subsections we show that for any $\varrho\in{\mathscr{P}}_{2}(\mathbb{R}^{n})$, $\gamma\in\Gamma(\rho)$, and $X_{0}\in{\mathcal{I}}^{\gamma}(\varrho)$ the reverse inequality holds,
$v(X_{0},\gamma)\geq v(X_{0}^{c},\gamma^{c}(\varrho)).$ (7.2)
Indeed, this together with Lemma 7.4 would imply that $v(X_{0},\gamma)=v(X_{0}^{c},\gamma^{c}(\varrho))$. Since the latter depends only on $\varrho$, we conclude for all possible bases and initial condition, the value function depends only on $\varrho$.
The above inequality is established in several steps. We first reduce it to the case of bounded cost functionals. The second step is to construct an approximating sequence of piecewise controls. We then transfer them to strong controls.
7.1 Reduction to Bounded Cost Functions
We start with the first reduction and suppress the dependence on $\gamma$. For positive constants $c,\ell>0$, set $F_{c}:=F\vee(-c)$, $G_{c}:=G\vee(-c)$ and
${\mathscr{A}}_{\ell}:=\{\alpha\in{\mathscr{A}}\,:\,|\alpha_{t}|\leq\ell\ \ \text{for all}\ t\in[0,T]\,\}.$
Let $J_{c}$ be the cost functional given by (3.1) with $(F,G)$ replaced with $(F_{c},G_{c})$ and define value functions by,
$v^{\ell}(X_{0}):=\inf_{\alpha\in{\mathscr{A}}_{\ell}}J(X_{0},\alpha),\qquad v^{\ell}_{c}(X_{0}):=\inf_{\alpha\in{\mathscr{A}}_{\ell}}J_{c}(X_{0},\alpha).$
We claim that
$\lim_{\ell\to\infty}v^{\ell}(X_{0})=v(X_{0}),\qquad\lim_{c\to\infty}v^{\ell}_{c}(X_{0})=v^{\ell}(X_{0}).$
To prove the first convergence, fix $a_{0}\in U$ and for any $\alpha\in{\mathscr{A}}$, set $\alpha^{\ell}_{t}:=\alpha_{t}$ if $|\alpha_{t}|\leq\ell$ and $\alpha^{\ell}_{t}:=a_{0}$ otherwise. Then, for large $\ell$, we have $\alpha^{\ell}\in{\mathscr{A}}_{\ell}$ and under the assumed regularity conditions, it is standard to show that $\lim_{\ell\to\infty}J(X_{0},\alpha^{\ell})=J(X_{0},\alpha)$. As $v^{\ell}$ is non-increasing in $\ell$, this implies that $\lim_{\ell\to\infty}v^{\ell}(X_{0})=v(X_{0})$. Also, for each $\alpha\in{\mathscr{A}}_{\ell}$, one can directly show that $\lim_{c\to\infty}J_{c}(X_{0},\alpha)=J(X_{0},\alpha)$. Since $v^{\ell}_{c}$ is non-increasing in $c$, these imply that $\lim_{c\to\infty}v^{\ell}_{c}(X_{0})=v^{\ell}(X_{0})$.
For a positive constant $m$ set $F^{m}_{c}:=F_{c}\wedge m$, $G_{c}^{m}:=G_{c}\wedge m$, and introduce the corresponding value functions by,
$v^{\ell}_{c,m}(X_{0}):=\inf_{\alpha\in{\mathscr{A}}_{\ell}}J^{m}_{c}(X_{0},\alpha),$
where $J^{m}_{c}$ is the cost functional defined using $(F^{m}_{c},G^{m}_{c})$.
Lemma 7.5.
Suppose that Assumptions 2.3 and 7.1 holds. Then, for every $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}_{0})$, and any $c,\ell$,
$\lim_{m\to\infty}v^{\ell}_{c,m}(X_{0})=v^{\ell}_{c}(X_{0}).$
Proof.
Without loss of generality assume $G\equiv 0$. We clearly have $v^{\ell}_{c,m}(X_{0})\leq v^{\ell}_{c}(X_{0})$ for any $m$. For the reverse inequality let $\alpha^{m}\in{\mathscr{A}}_{\ell}$ be a $1/m$-optimal control for $v^{\ell}_{c,m}(X_{0})$. Set
$\epsilon(m):=J_{c}(X_{0},\alpha^{m})-J^{m}_{c}(X_{0},\alpha^{m}).$
Then, $\epsilon(m)\geq 0$ and
$v^{\ell}_{c}(X_{0})\leq J_{c}(X_{0},\alpha^{m})=J^{m}_{c}(X_{0},\alpha^{m})+\epsilon(m)\leq v^{\ell}_{c,m}(X_{0})+1/m+\epsilon(m).$
Hence, it suffices to show $\lim_{m\to\infty}\epsilon(m)=0$. To establish this, we first note that the sequence $(X^{m})_{m\geq 1}$ is bounded in ${\mathscr{L}}^{2}$ due to boundedness of $\alpha^{m}\in{\mathscr{A}}_{\ell}$. In view of the quadratic growth Assumption 7.1 of $F,G$,
$0\leq\epsilon(m)$ $\leq\mathbb{E}\int_{0}^{T}|F_{c}(t,X^{m}_{\cdot\land t},\alpha^{m}_{t},{\mathcal{L}}^{m}_{t})-F_{c}^{m}(t,X^{m}_{\cdot\land t},\alpha^{m}_{t},{\mathcal{L}}^{m}_{t})|\,\mathrm{d}t$
$=\mathbb{E}\int_{0}^{T}|F_{c}(t,X^{m}_{\cdot\land t},\alpha^{m}_{t},{\mathcal{L}}^{m}_{t})|\chi_{\{|F_{c}(t,X^{m}_{\cdot\land t},\alpha^{m}_{t},{\mathcal{L}}^{m}_{t})|>m\}}\,\mathrm{d}t$
$\leq c_{*}\mathbb{E}\int_{0}^{T}[1+\|X^{m}\|_{t}^{2}+|\alpha^{m}|^{2}+\mathfrak{s}({\mathcal{L}}^{m}_{t})]\chi_{\{c_{*}[1+\|X^{m}\|_{t}^{2}+|\alpha^{m}|^{2}+\mathfrak{s}({\mathcal{L}}^{m}_{t})]>m\}}\,\mathrm{d}t,$
where we set $X^{m}:=X^{X_{0},\alpha^{m}}$, ${\mathcal{L}}^{m}:={\mathcal{L}}^{\alpha^{m}}$. By the boundedness of $\alpha^{m}$, the ${\mathscr{L}}^{2}$ boundedness of $(X^{m})_{m\geq 1}$, and basic properties of conditional expectations to deal with the ${\mathcal{L}}^{m}$-terms, it suffices to show uniform integrability of $(\|X^{m}\|_{T}^{2})_{m\geq 1}$,
$\lim_{K\to\infty}\ \sup_{m\geq 1}\ \mathbb{E}[\|X^{m}\|_{T}^{2}\chi_{\{\|X^{m}\|_{T}^{2}>K\}}]=0.$
Using the assumptions on the SDE coefficients in Assumption 7.1 and [42, Theorem 2.11], for any fixed $\beta\in{\mathscr{A}}_{\ell}$ and $\delta>0$, there exists a constant $c$ such that
$\mathbb{E}[\|X^{m}-X^{X_{0},\beta}\|_{T}^{2+\delta}]\leq c(1+\mathbb{E}\int_{0}^{T}\|X^{m}-X^{X_{0},\beta}\|_{t}^{2+\delta}\,\mathrm{d}t),$
and we conclude by Grönwall’s inequality that $\sup_{m\geq 1}\mathbb{E}[\|X^{m}-X^{X_{0},\beta}\|_{T}^{2+\delta}]\leq ce^{cT}$. Set $A(m,K):=\{\|X^{m}\|_{T}^{2}>K\}$. Clearly, by ${\mathscr{L}}^{2}$-boundedness of $(X^{m})_{m\geq 1}$,
$\sup_{m\geq 1}\mathbb{P}(A(m,K))\leq\frac{\sup_{m\geq 1}\mathbb{E}[\|X^{m}\|_{T}^{2}]}{K}\longrightarrow 0\quad\text{as}\ K\to\infty.$
This implies that, for any fixed $\beta\in{\mathscr{A}}_{\ell}$, using Hölder’s inequality with $p:=(2+\delta)/2$ and $q:=(2+\delta)/\delta$,
$\mathbb{E}[\|X^{m}\|_{T}^{2}\chi_{A(m,K)}]$ $\leq 2\,\mathbb{E}([\|X^{m}-X^{X_{0},\beta}\|_{T}^{2}+\|X^{X_{0},\beta}\|_{T}^{2})\chi_{A(m,K)}]$
$\leq 2\,\mathbb{E}[\|X^{m}-X^{X_{0},\beta}\|_{T}^{2+\delta}]^{1/p}\mathbb{P}(A(m,K))^{1/q}+2\,\mathbb{E}[\|X^{X_{0},\beta}\|_{T}^{2}\chi_{A(m,K)}]$
$\leq 2\,(ce^{cT})^{1/p}\mathbb{P}(A(m,K))^{1/q}+2\,\mathbb{E}[\|X^{X_{0},\beta}\|_{T}^{2}\chi_{A(m,K)}].$
This converges to zero as $K\to\infty$, uniformly in $m$, completing the proof. ∎
Now suppose that Theorem 7.2 holds for all $F$ and $G$ that satisfy not only Assumption 7.1 but are also bounded. This would imply that for every $\ell,c,m$, the invariance principle holds for the MFC with functions $(F^{m}_{c},G^{m}_{c})$ and with ${\mathscr{A}}_{\ell}$. Then, for any $\gamma,\gamma^{\prime}\in\Gamma(\varrho)$, $X_{0}\in{\mathcal{I}}^{\gamma}(\varrho)$ and $X_{0}\in{\mathcal{I}}^{\gamma}(\varrho)$, $X^{\prime}_{0}\in{\mathcal{I}}^{\gamma^{\prime}}(\varrho)$,
$v^{\ell}_{c,m}(X_{0},\gamma)=v^{\ell}_{c,m}(X_{0}^{\prime},\gamma^{\prime}).$
In view of the above lemma and the discussion preceding it, this implies that $v(X_{0},\gamma)=v(X_{0}^{\prime},\gamma^{\prime})$, proving the invariance principle for the original cost functions $(F,G)$. Hence, without loss of generality we may assume that $F$, $G$ are bounded and by Assumption 7.1, there are uniformly continuous with a bounded continuous modulus of continuity $m$.
7.2 Delayed Start
This section introduces a technique presented in [25]. We now fix a stochastic basis $\gamma$, an initial condition $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}^{\gamma}_{0})$, and suppress the dependence on $\gamma$ in this section. For $\alpha\in{\mathscr{A}}$ and $\epsilon>0$, we construct an approximating state process $X^{\alpha,\epsilon}$ and its cost. We first set
$X^{\alpha,\epsilon}_{t}=X_{0},\qquad\forall\,t\in[0,\epsilon].$
For $t\in[\epsilon,T]$, we let $X^{\alpha,\epsilon}$ be the unique strong solution of the following equation,
$X^{\alpha,\epsilon}_{t}=X_{0}+\int_{\epsilon}^{t}$ $b(s,X^{\alpha,\epsilon}_{\cdot\wedge s},\alpha_{s})\,\mathrm{d}s+\int_{\epsilon}^{t}\boldsymbol{\sigma}(s,X^{\alpha,\epsilon}_{\cdot\wedge s},\alpha_{s})\,\mathrm{d}{\mathscr{W}}_{s}$
[…]
Using ${\mathcal{L}}^{\alpha,\epsilon}_{t}:={\mathcal{L}}((X^{\alpha,\epsilon}_{\cdot\wedge t},\alpha_{t})\,|\,{\mathscr{G}})$, we define the approximating cost functional by,
$J^{\epsilon}(X_{0},\alpha):=\mathbb{E}\left[\int_{0}^{T}F(t,X^{\alpha,\epsilon}_{\cdot\wedge t},\alpha_{t},{\mathcal{L}}^{\alpha,\epsilon}_{t})\ \mathrm{d}t+G(X^{\alpha,\epsilon}_{\cdot\wedge T},{\mathcal{L}}^{\alpha,\epsilon}_{T})\right].$ (7.3)
Recall $\|\alpha\|_{2,T}$ of (2.1).
Lemma 7.6.
Consider sequences $\alpha^{k}\in{\mathscr{A}}$, $\epsilon_{k}>0$. Then,
$\lim_{k\to\infty}\ (\epsilon_{k}+\|\alpha^{k}\|_{2,\epsilon_{k}})=0\quad\Rightarrow\quad\lim_{k\to\infty}\ |J(X_{0},\alpha^{k})-J^{\epsilon_{k}}(X_{0},\alpha^{k})|=0.$
Proof.
We first fix $\alpha\in{\mathscr{A}}$, $\epsilon>0$, and we suppress the dependence on $\mathbb{E}|X_{0}|^{2}$, and on the constants in Assumptions 2.3 and 7.1. To simplify the notation, we set
$X^{k}:=X^{\alpha^{k}},\quad Y^{k}:=X^{\alpha^{k},\epsilon_{k}},\quad{\mathcal{L}}^{k}_{t}:={\mathcal{L}}^{\alpha^{k}}_{t},\quad\widehat{\mathcal{L}}^{k}_{t}:={\mathcal{L}}^{\alpha^{k},\epsilon_{k}}_{t}.$
By Assumption 2.3, there is a constant $c_{1}$ such that,
$\mathbb{E}[\|X^{\alpha}-X^{\alpha,\epsilon}\|^{2}_{\epsilon}]=\mathbb{E}[\sup_{s\in[0,\epsilon]}|X^{\alpha}_{t}-X_{0}|^{2}]\leq c_{1}(\epsilon+\|\alpha\|^{2}_{2,\epsilon}).$
We then use the Burkholder-Davis-Gundy inequality, [42, Corollary 2.12], the Lipschitz property of the coefficients $(b,\boldsymbol{\sigma},\boldsymbol{\lambda})$, and the Grönwall’s inequality to arrive at
$\mathbb{E}[\sup_{t\in[\epsilon,T]}|X^{\alpha}_{t}-X^{\alpha,\epsilon}_{t}|^{2}]\ \leq\ c_{2}\mathbb{E}[|X^{\alpha}_{\epsilon}-X^{\alpha,\epsilon}_{\epsilon}|^{2}]=c_{2}\mathbb{E}[|X^{\alpha}_{\epsilon}-X_{0}|^{2}],$
for some constant $c_{2}$. Hence, there is $c_{3}$ such that
$\mathbb{E}[\|X^{\alpha}-X^{\alpha,\epsilon}\|^{2}_{T}]\leq\mathbb{E}[\|X^{\alpha}-X^{\alpha,\epsilon}\|^{2}_{\epsilon}]+\mathbb{E}[\sup_{t\in[\epsilon,T]}|X^{\alpha}_{t}-X^{\alpha,\epsilon}_{t}|^{2}]\leq\ c_{3}(\epsilon+\|\alpha\|^{2}_{2,\epsilon}).$
Since $\mathbb{E}[\sup_{t\in[0,T]}\mathfrak{m}_{2}({\mathcal{L}}^{\alpha}_{t}\,,\,{\mathcal{L}}^{\alpha,\epsilon}_{t})^{2}]\leq\mathbb{E}[\|X^{\alpha}-X^{\alpha,\epsilon}\|^{2}_{T}]$ it follows that $\lim_{k\to\infty}\mathbb{E}[\zeta_{k}]=0$, where
$\zeta_{k}:=\|X^{k}-Y^{k}\|^{2}_{T}+\sup_{t\in[0,T]}\mathfrak{m}_{2}({\mathcal{L}}^{k}_{t}\,,\,\hat{\mathcal{L}}^{k}_{t})^{2}.$
Hence, $\zeta_{k}$ converges to zero in probability. Moreover, by the previous subsection,
$|F(t,X^{k}_{\cdot\land t},\alpha^{k}_{t},{\mathcal{L}}^{k}_{t})-F(t,Y^{k}_{\cdot\land t},\alpha^{k}_{t},\widehat{\mathcal{L}}^{k}_{t})|+|G(X^{k}_{\cdot\land T},{\mathcal{L}}^{k}_{T})-G(Y^{k}_{\cdot\land T},\widehat{\mathcal{L}}^{k}_{T})|\leq m(\zeta_{k}),$
for every $t\in[0,T]$, where $m$ is the modulus of continuity of $F$ and $G$. Consequently, the following sequence of random variables
$j_{k}:=\int_{0}^{T}[F(t,X^{k}_{\cdot\land t},\alpha^{k}_{t},{\mathcal{L}}^{k}_{t})-F(t,Y^{k}_{\cdot\land t},\alpha^{k}_{t},\widehat{\mathcal{L}}^{k}_{t})]\,\mathrm{d}t+G(X^{k}_{\cdot\land T},{\mathcal{L}}^{k}_{T})-G(Y^{k}_{\cdot\land T},\widehat{\mathcal{L}}^{k}_{T}),$
also converge to zero in probability. Additionally, by the reduction obtained in the previous subsection, $F$ and $G$ are bounded. Therefore, $|j_{k}|\leq 2\|F\|_{\infty}T+2\|G\|_{\infty}$. Then, by classical arguments we arrive at
$\lim_{k\to\infty}|J(X_{0},\alpha^{k})-J^{,\epsilon_{k}}(X_{0},\alpha^{k})|=\lim_{k\to\infty}\mathbb{E}[|j_{k}|]=0.$
7.3 Simple Controls
Definition 7.7.
A control process $\alpha\in{\mathscr{A}}(\gamma)$ is called simple, if
$\alpha_{t}=a_{0}\chi_{[0,t_{1}]}(t)+\sum_{i=1}^{m}\alpha_{i}\chi_{(t_{i},t_{i+1}]}(t),\qquad t\in[0,T],$ (7.4)
for some $t_{0}=0<t_{1}=:t_{1}(\alpha)<..<t_{m+1}=T$, ${\mathscr{F}}_{t_{i}}^{\gamma}$-measurable $\alpha_{i}:\Omega^{\gamma}\mapsto U$ and $a_{0}\in U$.
Let ${\mathscr{A}}_{s}(\gamma)$ be the set of all simple controls, and for any $a_{0}\in U$, ${\mathscr{A}}_{s}(\gamma,a_{0})$ is the set of all simple controls that start with $a_{0}$.
Lemma 7.8.
Let $\Xi_{T}^{\gamma,\epsilon}$ be as in (7.1). For any bases $\gamma,\tilde{\gamma}$, $\epsilon>0$, $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}^{\gamma}_{0})$, $\alpha\in{\mathscr{A}}_{s}(\gamma)$, $\tilde{X}_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}^{\tilde{\gamma}}_{0})$, and $\tilde{\alpha}\in{\mathscr{A}}_{s}(\tilde{\gamma})$ with $t_{1}(\alpha)=t_{1}(\tilde{\alpha})=\epsilon$,
${\mathcal{L}}^{\gamma}(X_{0},\alpha,\Xi^{\gamma,\epsilon}_{T})={\mathcal{L}}^{\tilde{\gamma}}(\tilde{X}_{0},\tilde{\alpha},\Xi^{\tilde{\gamma},\epsilon}_{T})\quad\Rightarrow\quad J_{g}^{\gamma,\epsilon}(X_{0},\alpha)=J_{g}^{\tilde{\gamma},\epsilon}(\tilde{X}_{0},\tilde{\alpha}).$
Proof.
Consider the measurable space ${\mathscr{Y}}:=\mathbb{R}^{n}\times U^{m}\times{\mathscr{C}}_{\ell}\times{\mathscr{C}}_{d}\times M^{2}$. Then, ${\mathcal{L}}^{\gamma}(X_{0},\alpha,\Xi^{\gamma,\epsilon}_{T})$ is the measure induced by the map, $\omega\in\Omega^{\gamma}\mapsto(X_{0}(\omega),\alpha_{1}(\omega),..,\alpha_{m}(\omega),\Xi^{\gamma,\epsilon}_{T}(\omega))\in{\mathscr{Y}}$. Since the state process $X^{\alpha,\epsilon}$ is a function of the differentials of $({\mathscr{W}}_{t},{\mathscr{N}}_{t})$ for $t\in[\epsilon,T]$, the independent increment property of these processes imply that law of $X^{\alpha,\epsilon}$ depends only on ${\mathcal{L}}^{\gamma}(X_{0},\alpha,\Xi^{\gamma,\epsilon}_{T})$. Then, the statement follows directly from the definition of $J_{g}^{\gamma,\epsilon}$. ∎
Lemma 7.9.
Suppose Assumption 7.1 hold. Then, for any basis $\gamma$, $a_{0}\in U$, initial condition $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}^{\gamma}_{0})$ and control $\alpha\in{\mathscr{A}}(\gamma)$, there exists $\alpha^{k}\in{\mathscr{A}}_{s}(\gamma,a_{0})$ such that with $\epsilon_{k}:=t_{1}(\alpha^{k})$,
$\lim_{k\to\infty}\epsilon_{k}=0,\qquad\text{and}\qquad\lim_{k\to\infty}J^{\gamma,\epsilon_{k}}(X_{0},\alpha^{k})=J^{\gamma}(X_{0},\alpha).$ (7.5)
Proof.
In the absence of jumps, this result is proved in [26, Lemma 2], and we follow its proof closely. We first fix $\alpha\in{\mathscr{A}}(\gamma)$, and construct a sequence of simple controls $\alpha^{k}$ approximating $\alpha$ in the ${\mathscr{L}}^{2}(\mathrm{d}\mathbb{P}\otimes\mathrm{d}t)$-norm, [50, Lemma 4.4]. Then, the same estimates as in Lemma 7.6 show that $\lim_{k\to\infty}\mathbb{E}[\|X^{\alpha^{k},\epsilon^{k}}-X^{\alpha}\|_{T}^{2}]=0$. Following the proof of the aforementioned lemma, we use the continuity and boundedness of the functions $F,G$ to conclude that the cost functionals converge. ∎
7.4 Strong Controls
Definition 7.10.
A control process $\alpha\in{\mathscr{A}}(\gamma)$ is called strong, if
$\alpha_{t}(\omega)=\alpha_{t}(X_{0}(\omega),\Xi_{t}^{\gamma}(\omega)),\qquad t\in[0,T],\ \omega\in\Omega^{\gamma},$
where $\Xi_{t}^{\gamma}$ is defined in (7.1).
We let ${\mathscr{A}}_{st}(\gamma)$ be the set of all strong controls. It is clear that there is a one-to-one connection between ${\mathscr{A}}_{st}(\gamma)$ and ${\mathscr{A}}_{st}(\gamma^{c}(\varrho))$. Indeed,
$\forall\alpha\in{\mathscr{A}}_{st}(\gamma),\ \ \exists\alpha^{c}\in{\mathscr{A}}_{st}(\gamma^{c}(\varrho))\ \ \text{such that}\ \ J^{\gamma}(X_{0},\alpha)=J^{\gamma^{c}(\varrho)}(X_{0}^{c},\alpha^{c}).$ (7.6)
We next construct a strong control for each simple control in $\gamma$, by utilizing a classical stochastic transfer lemma, see for instance [40, Theorem 6.10].
Lemma 7.11 (Stochastic transfer theorem).
Suppose that $(S,{\mathcal{S}})$ is a measurable space and $(T,{\mathcal{T}})$ is a Borel space. Let $Y$ and $Z$ be random variables on a probability space $(\Omega,{\mathscr{F}},\mathbb{P})$ taking values in $S$ and $T$, respectively. Then, for any uniform random variable $U\in[0,1]$ independent of $X$, there exists a measurable function $\psi:[0,1]\times S\to T$ such that ${\mathcal{L}}^{\mathbb{P}}(Y,\psi(U,Y))={\mathcal{L}}^{\mathbb{P}}(Y,Z)$.
The following is the main technical step of the proof, and it is inspired by [26, Lemma 3].
Proposition 7.12.
For $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}_{0}^{\gamma})$, $a_{0}\in U$, and a simple control $\alpha\in{\mathscr{A}}_{s}(\gamma,a_{0})$, there is a strong control $\alpha^{st}\in{\mathscr{A}}_{st}(\gamma)\cap{\mathscr{A}}_{s}(\gamma,a_{0})$ such that with $\epsilon=t_{1}(\alpha)$ the followings hold,
$t_{1}(\alpha^{st})=\epsilon,\qquad{\text{and}}\qquad J^{\gamma,\epsilon}(X_{0},\alpha)=J^{\gamma,\epsilon}(X_{0},\alpha^{st}).$
Proof.
We fix the basis $\gamma$, $X_{0}\in{\mathcal{I}}^{\gamma}(\varrho)$, $a_{0}\in U$, and $\alpha\in{\mathscr{A}}_{s}(\gamma,a_{0})$. With $\alpha$ as in (7.4), set $\epsilon:=t_{1}(\alpha)$ and
$\Xi_{i}:=\Xi^{\gamma,\epsilon}_{t_{i}}\in{\mathscr{C}}_{\ell}\times{\mathscr{C}}_{d}\times M^{2}=:{\mathscr{X}},\qquad i=1,..,m,$
where $\Xi^{\gamma,\epsilon}_{t}$ is as in (7.1). Then, $(X_{0},\Xi_{1},..,\Xi_{m})$ is independent of $W^{\gamma}_{\epsilon}$.
It is classical that there is a measurable map $\varphi:\mathbb{R}^{d}\mapsto[0,1]^{m}$ such that components of $\varphi(W^{\gamma}_{\epsilon})=:\xi=:(\xi_{1},..,\xi_{m})$ are independent uniform variables on $[0,1]$. Additionally, $(X_{0},\Xi_{1},..,\Xi_{m})$ and $\xi$ are independent.
Step 1. By the stochastic transfer lemma with $Y=X_{0}$, $Z=\alpha_{1}$, and $U=\xi_{1}$, there exists a measurable $\psi_{1}:[0,1]\times\mathbb{R}^{n}\mapsto U$ such that ${\mathcal{L}}(X_{0},\psi(\xi_{1},X_{0}))={\mathcal{L}}(X_{0},\alpha_{1})$. Then, for $i=2,..,m$, we apply the transfer lemma with $Y=(X_{0},\alpha_{1},..,\alpha_{i-1},\Xi_{i})$, $Z=\alpha_{i}$, and $U=\xi_{i}$ to construct a measurable map $\psi_{i}:[0,1]\times U^{i-1}\times\mathbb{R}^{n}\times{\mathscr{X}}\mapsto U$ such that
${\mathcal{L}}(X_{0},\alpha_{1},..,\alpha_{i-1},\Xi_{i},\psi_{i}(\xi_{i},X_{0},\alpha_{1},..,\alpha_{i-1},\Xi_{i}))={\mathcal{L}}(X_{0},\alpha_{1},..,\alpha_{i-1},\Xi_{i},\alpha_{i}).$
Step 2. We set $\alpha^{st}_{1}:=\psi_{1}(\xi_{1},X_{0})$. Then, recursively we define
$\alpha^{st}_{i}:=\psi_{i}(\xi_{i},X_{0},\alpha^{st}_{1},..,\alpha^{st}_{i-1},\Xi_{i}),\quad i=2,..,m.$
Using this definition and the independence of increments, one may recursively show that
${\mathcal{L}}(X_{0},\alpha_{1},..,\alpha_{i},\Xi_{i})={\mathcal{L}}(X_{0},\alpha^{st}_{1},..,\alpha^{st}_{i},\Xi_{i}),\qquad i=1,..,m.$
Step 3. Since $\Theta^{st}:=(X_{0},\alpha^{st}_{1},..,\alpha^{st}_{m},\Xi_{m})$ and $\Theta:=(X_{0},\alpha_{1},..,\alpha_{m},\Xi_{m})$ are ${\mathscr{F}}^{\gamma}_{t_{m}}$ measurable, they are both independent of the increment $\Xi^{\gamma,\epsilon}_{T}-\Xi_{m}=\Xi^{\gamma,\epsilon}_{T}-\Xi^{\gamma,\epsilon}_{t_{m}}$. Also, by the previous step ${\mathcal{L}}(\Theta^{st})={\mathcal{L}}(\Theta)$, hence ${\mathcal{L}}(\Theta^{st},\Xi^{\gamma,\epsilon}-\Xi_{m})={\mathcal{L}}(\Theta,\Xi^{\gamma,\epsilon}-\Xi_{m})$. Therefore,
${\mathcal{L}}(X_{0},\alpha^{st}_{1},..,\alpha^{st}_{m},\Xi_{T}^{\gamma,\epsilon})={\mathcal{L}}(X_{0},\alpha_{1},..,\alpha_{m},\Xi_{T}^{\gamma,\epsilon}).$
Step 4. We define a simple, strong control process by,
$\alpha^{st}_{t}:=a_{0}\chi_{[0,t_{1}]}(t)+\sum_{i=1}^{m}\alpha^{st}_{i}\chi_{(t_{i},t_{i+1}]}(t),\qquad t\in[0,T].$
In view of Step 2, $\alpha^{st}\in{\mathscr{A}}_{st}(\gamma)\cap{\mathscr{A}}_{s}(\gamma,a_{0})$. By the previous step,
${\mathcal{L}}(X_{0},\alpha^{st},\Xi_{T}^{\gamma,\epsilon})={\mathcal{L}}(X_{0},\alpha,\Xi_{T}^{\gamma,\epsilon}).$
Then, by Lemma 7.8, $J^{\gamma,\epsilon}(X_{0},\alpha)=J^{\gamma,\epsilon}(X_{0},\alpha^{st})$, and $\|\alpha\|_{2,T}=\|\alpha^{st}\|_{2,T}$. ∎
7.5 Proof of Theorem 7.2
For $\delta>0$, let $\alpha^{\delta}\in{\mathscr{A}}(\gamma)$ be a $\delta$-optimizer of the MFC problem on $\gamma$, i.e.,
$v(X_{0},\gamma)\geq J^{\gamma}(X_{0},\alpha^{\delta})-\delta.$
We fix $a_{0}\in U$. By Lemma 7.9, there are $\alpha^{k}\in{\mathscr{A}}_{s}(\gamma,a_{0})$ such that with $\epsilon_{k}=t_{1}(\alpha^{k})\downarrow 0$,
$\lim_{k\to\infty}\ J^{\gamma,\epsilon_{k}}(X_{0},\alpha^{k})=J^{\gamma}(X_{0},\alpha^{\delta}).$
By Proposition 7.12, there are $\alpha^{k,st}\in{\mathscr{A}}_{st}(\gamma)\cap{\mathscr{A}}_{s}(\gamma,a_{0})$ satisfying,
$J^{\gamma,\epsilon_{k}}(X_{0},\alpha^{k})=J^{\gamma,\epsilon_{k}}(X_{0},\alpha^{k,st})\quad\text{and}\quad t_{1}(\alpha^{k,st})=\epsilon_{k}.$
Since $\alpha^{k,st}\in{\mathscr{A}}_{s}(\gamma,a_{0})$, $\alpha^{k,st}_{t}=a_{0}$ for every $t\in[0,\epsilon_{k}]$ and $\|\alpha^{k,st}\|^{2}_{2,\epsilon_{k}}=|a_{0}|^{2}\epsilon_{k}$. Hence, by Lemma 7.6,
$\lim_{k\to\infty}\ |J^{\gamma}(X_{0},\alpha^{k,st})-J^{\gamma,\epsilon_{k}}(X_{0},\alpha^{k,st})|=0.$
By (7.6), there are $\alpha^{k,st,c}\in{\mathscr{A}}(\gamma^{c}(\varrho))$ so that
$J^{\gamma}(X_{0},\alpha^{k,st})=J^{\gamma^{c}(\varrho)}(X_{0}^{c},\alpha^{k,st,c}).$
Using these we choose $\alpha^{\delta,1}\in{\mathscr{A}}_{s}(\gamma)$, $\alpha^{\delta,2}\in{\mathscr{A}}_{st}(\gamma)$, $\alpha^{\delta,c}\in{\mathscr{A}}(\gamma^{c}(\varrho))$, and $\epsilon_{*}>0$ satisfying,
$v(X_{0},\gamma)$ $\geq J^{\gamma}(X_{0},\alpha^{\delta})-\delta\geq J^{\gamma,\epsilon_{*}}(X_{0},\alpha^{\delta,1})-2\delta=J^{\gamma,\epsilon_{*}}(X_{0},\alpha^{\delta,2})-2\delta$
$\geq J^{\gamma}(X_{0},\alpha^{\delta,2})-3\delta=J^{\gamma^{c}(\varrho)}(X_{0}^{c},\alpha^{\delta,c})-3\delta\geq v(X_{0}^{c},\gamma^{c}(\varrho))-3\delta.$
As $\delta>0$ is arbitrary, we conclude that $v(X_{0},\gamma)\geq v(X_{0}^{c},\gamma^{c}(\varrho))$. This is exactly (7.2) which, as already argued earlier, implies the invariance principle.
$\Box$
Appendix A Poisson Point Processes
In this subsection, for the convenience of the reader we recall standard notations and terminology from the classical textbook of Ikeda & Watanabe [39] on stochastic processes. See also [2] for details on stochastic differential equations with jumps and the construction of canonical spaces for random measures.
Let $(E,{\mathscr{E}})=(\mathbb{R}^{k}\setminus\{0\},{\mathscr{B}}(\mathbb{R}^{k}\setminus\{0\}))$, and a counting measure on $[0,\infty)\times E$ has the form
$\pi=\sum_{i\in{\mathcal{I}}}\delta_{\{(t_{i},p_{i})\}},$
[…]
A ($\sigma$-finite) integer-valued random measure is any $(M,{\mathscr{M}})$-valued random variable. A Poisson random measure on $[0,\infty)\times E$ is a random measure that satisfies
  1. (i)
    For $B\in{\mathscr{B}}[0,\infty)\otimes{\mathscr{E}}$, $\pi(B)$ is Poisson with the intensity measure $\lambda(B)=\mathbb{E}[\pi(B)]$;
  2. (ii)
    For disjoint sets $B_{1},..,B_{n}\in{\mathscr{B}}[0,\infty)\otimes{\mathscr{E}}$, the random variables $\pi(B_{1}),..,\pi(B_{n})$ are mutually independent.
A function $p:D(p)\subset(0,\infty)\to E$ is called a point function on $E$, if its domain $D(p)$ is a countable and $\{s\in D(p)\cap(0,t]:p(s)\in S\}$ finite for every $t>0$ and a compact $S\subset E$. For every point function on $E$ we associate a counting measure $N_{p}$ on $[0,\infty)\times E$ by,
$N_{p}(\mathrm{d}t,\mathrm{d}\zeta)=\sum_{s\in D_{p}}\delta_{\{(s,p(s))\}}(\mathrm{d}t,\mathrm{d}\zeta).$
Note that $N_{p}(\{0\}\times E)=0$. We equip the space of all point functions with the $\sigma$-algebra which is generated by the family of maps $p\mapsto N_{p}((0,t]\times S)\in[0,\infty]$, for $t>0$, $S\in{\mathscr{E}}$.
A point process is a randomized point function $p(\omega,t)$, measurable in the sense introduced before, and it is called a Poisson point process if the associated random counting measure $N_{p}$ is a Poisson random measure. Finally, it is called stationary, if the intensity measure of $N_{p}$ is given by the product of Lebesgue measure with some $\sigma$-finite measure $n(\mathrm{d}\zeta)$ on $(E,{\mathscr{E}})$:
$\hat{N}_{p}(\mathrm{d}t,\mathrm{d}\zeta)=\mathrm{d}t\otimes n(\mathrm{d}\zeta).$
The deterministic measure $\hat{N}_{p}$ is the compensator or dual predictable projection of the Poisson random measure $N_{p}$. The random measure $\widetilde{N}_{p}:=N_{p}-\hat{N}_{p}$ is called the compensated Poisson random measure. With these definitions, $\sigma$-finite integer-valued random measures are in one-to-one correspondence with point processes.
On a filtered probability space $(\Omega,{\mathscr{F}},\mathbb{F},\mathbb{P})$, a Poisson point process $p=p(\omega,t)$ on $\Omega$ is called an $\mathbb{F}$-Poisson point process if for every $t,h>0$, and $S\in{\mathscr{E}}$, $N_{p}((0,t]\times S)$ is ${\mathscr{F}}_{t}$-measurable, and the increments $(N_{p}((0,t+h]\times S)-N_{p}((0,t]\times S))$ are independent of ${\mathscr{F}}_{t}$. For such processes, the stochastic integral
$\int_{0}^{t}\int_{E}f(\omega,s,\zeta)\,\widetilde{N}_{p(\omega)}(\mathrm{d}s,\mathrm{d}\zeta)$
is defined for any real function $f$ on $\Omega\times[0,\infty)\times E$ that is ${\mathscr{P}}\otimes{\mathscr{E}}$-measurable, where ${\mathscr{P}}$ is the predictable $\sigma$-field on $\Omega\times[0,\infty)$, and which satisfies
$\mathbb{P}\left(\int_{0}^{t}\int_{E}|f(\omega,s,\zeta)|^{2}\,n(\mathrm{d}\zeta)\,\mathrm{d}s<\infty\right)=1,\quad t>0.$
Moreover, the identity
$\int_{0}^{t}\int_{E}f(s,\zeta)\,\widetilde{N}_{p}(\mathrm{d}s,\mathrm{d}\zeta)=\int_{0}^{t}\int_{E}f(s,\zeta)\,N_{p}(\mathrm{d}s,\mathrm{d}\zeta)-\int_{0}^{t}\int_{E}f(s,\zeta)\,n(\mathrm{d}\zeta)\,\mathrm{d}s,$
with the right side being standard Lebesgue integrals, only holds when the terms on the right side separately converge. In general, the left side is defined as an (localized) ${\mathscr{L}}^{2}$-limit. For details we refer to Chapter II.3 [39] and [42] which present a more general integration theory with respect to random compensators.
Appendix B Discounted Infinite Horizon
We consider the discounted infinite horizon problems with a fixed discount factor $\beta>0$.
Definition B.1 ($\beta$-admissible controls).
For a probability basis $\gamma$, the set of $\beta$-admissible controls ${\mathscr{A}}_{\beta}(\gamma)$ is the set of all $\mathbb{F}^{\gamma}$-progressively measurable processes $\alpha:\Omega^{\gamma}\times[0,\infty)\to U$ satisfying
$\mathbb{E}^{\gamma}\int_{0}^{\infty}e^{-\beta t}[\|X^{\alpha}\|^{2}_{t}+|\alpha_{t}|^{2}]\,\mathrm{d}t<\infty,$ (B.1)
for any initial condition $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}^{\gamma}_{0})$. Here $X^{\alpha}$ is the unique strong solution to SDE$(X_{0},\alpha)$.
The discounted infinite horizon MFC problem is to minimize
$J_{\beta}^{\gamma}(X_{0},\alpha):=\mathbb{E}^{\gamma}\int_{0}^{\infty}e^{-\beta t}F(t,X^{\alpha}_{\cdot\land t},\alpha_{t},{\mathcal{L}}^{\alpha}_{t})\,\mathrm{d}t,$ (B.2)
over all $\alpha\in{\mathscr{A}}_{\beta}(\gamma)$, where as before ${\mathcal{L}}_{t}^{\alpha}:={\mathcal{L}}^{\gamma}((X^{\alpha}_{\cdot\land t},\alpha_{t})\,|\,{\mathscr{G}}^{\gamma})$.
For a given $\varrho\in{\mathscr{P}}_{2}(\mathbb{R}^{n})$, ${\mathcal{I}}_{\beta}^{\gamma}(\varrho)$ is the set of all ${\mathscr{F}}_{0}^{\gamma}$-measurable $X_{0}$ such that ${\mathcal{L}}^{\gamma}(X_{0})=\varrho$, and $\Gamma_{\beta}(\varrho)$ is the set of all $\gamma$ such that both ${\mathscr{A}}_{\beta}(\gamma)$ and ${\mathcal{I}}_{\beta}(\varrho)$ are non-empty.
Remark B.2.
In many applications, the controlled diffusion follows $\mathrm{d}X^{\alpha}_{t}=\alpha_{t}\,\mathrm{d}t+\sigma\,\mathrm{d}W_{t}$. In this case, $\alpha\equiv 0$ always belong to ${\mathscr{A}}_{\beta}(\gamma)$ and the additional requirement about the existence of a control is always satisfied.
The following result is obtained by following the proof of Theorem 7.2 mutatis mutandis.
Theorem B.3.
Suppose that Assumptions 2.3, 7.1 hold. Then, for any $\varrho\in{\mathscr{P}}_{2}(\mathbb{R}^{n})$,
$v_{\beta}(X_{0},\gamma):=\inf_{\alpha\in{\mathscr{A}}_{\beta}(\gamma)}\,\,J_{\beta}^{\gamma}(X_{0},\alpha)=\inf_{\gamma\in\Gamma_{\beta}(\varrho)}\inf_{X_{0}\in{\mathcal{I}}_{\beta}^{\gamma}(\varrho)}\inf_{\alpha\in{\mathscr{A}}_{\beta}(\gamma)}J_{\beta}^{\gamma}(X_{0},\alpha)=:V_{\beta}(\varrho),$
for all $\gamma\in\Gamma_{\beta}(\varrho)$ and $X_{0}\in{\mathcal{I}}_{\beta}^{\gamma}(\varrho)$.
We continue by defining the MFG in this context. Let $f$ be as in (1.4), and set
$J_{g,\beta}^{\gamma}(X_{0},\alpha,\nu^{*}):=\mathbb{E}^{\gamma}\int_{0}^{\infty}e^{-\beta t}f(t,X^{\alpha}_{\cdot\land t},\alpha_{t},\nu^{*}_{t})\,\mathrm{d}t,$
Definition B.4.
For a basis $\gamma$ and an initial condition $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}^{\gamma}_{0})$, a pair $(\alpha^{*},\nu^{*})$ is an discounted MFG of controls Nash equilibrium on $\gamma$ if the followings hold:
  1. (i)
    $\alpha^{*}\in\mathcal{A}_{\beta}(\gamma)$ satisfies $J_{g,\beta}^{\gamma}(X_{0},\alpha^{*},\nu^{*})=\inf_{\alpha\in{\mathscr{A}}_{\beta}(\gamma)}J_{g,\beta}^{\gamma}(X_{0},\alpha,\nu^{*})$,
  2. (ii)
    $\nu^{*}_{t}={\mathcal{L}}^{\gamma}((X^{\alpha^{*}}_{\cdot\land t},\alpha^{*}_{t})\,|\,{\mathscr{G}}^{\gamma}\,)$ $\mathbb{P}^{\gamma}$-a.s. for all $t\geq 0$.
A straightforward modification of the proof of Theorem 3.6, using the integrability condition (B.1) and Theorem B.3, shows the following.
Theorem B.5.
Suppose that Assumptions 2.3, 3.5 hold, and the MFC problem is invariant under probability bases as in Theorem B.3. Let $\gamma$ be a probabilistic basis with non-empty ${\mathscr{A}}_{\beta}(\gamma)$ and $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}^{\gamma}_{0})$. If $\alpha^{*}\in{\mathscr{A}}_{\beta}(\gamma)$ is a minimizer of the MFC problem (B.2), then $(\alpha^{*},\nu^{*})$ is a discounted MFG of controls Nash equilibrium for the potential game on the basis $\gamma$, where $\nu^{*}_{t}:={\mathcal{L}}^{\gamma}((X^{\alpha^{*}}_{\cdot\land t},\alpha^{*}_{t})\,|\,{\mathscr{G}}^{\gamma}\,)$.